### v2

# arXiv:2603.00989v2[cs.SE] 03 Mar 2026

Abstract
Large Language Models (LLMs) are increasingly used in software engineering to help developers generate, complete, translate, and fix code. These tools improve productivity and speed up the development process. However, their growing use has raised concerns about environmental sustainability. Most existing studies focus on the high energy consumption and carbon emissions of model training and inference. In contrast, much less attention has been given to the sustainability of the code generated by these models. The efficiency of generated code affects the long-term environmental impact of software systems. Inefficient code can increase CPU usage, memory consumption, execution time, and overall energy use during deployment and operation. As LLM-generated code becomes more common in real-world projects, even small inefficiencies can lead to high environmental costs over time. This paper examines existing research on the sustainability of code generated by LLMs. We conduct a systematic literature review to analyze selected primary studies and investigate the extent to which LLMs are capable of producing sustainable code. In addition, we examine how sustainability is defined and measured in this context, including the metrics and evaluation strategies used to assess energy efficiency and resource usage. We also explore whether techniques such as fine-tuning and prompt engineering influence the sustainability of generated code. Through a structured analysis of the selected studies, we categorize research efforts based on their methodological approaches, evaluation practices, and experimental settings. The findings indicate that research in this area remains relatively limited and fragmented, with no widely accepted framework for measuring or benchmarking the sustainability of LLM-generated code. These observations highlight the need for clearer definitions, standardized evaluation methods, and systematic research to support environmentally friendly AI-assisted software engineering.
I Introduction
Large Language Models (LLMs) are advanced AI systems trained on large corpora of data using transformer-based architectures [131]. These models are capable of understanding, reasoning over, and generating human-like language, enabling their use across a wide range of applications[150]. In recent years, LLMs have gained significant attention in the software engineering domain, where they support and automate various programming tasks. One of the most prominent applications is code generation, in which an LLM produces executable source code from natural language instructions. Closely related tasks include code completion (predicting missing or subsequent lines of code), code translation (converting source code between programming languages), bug fixing (detecting and correcting programming errors), and documentation generation (producing natural language explanations of code functionality). These capabilities are now embedded in widely used AI-assisted development tools, such as GitHub Copilot [43], CodeGeeX [144], and Amazon CodeWhisperer [8], which leverage state-of-the-art code LLMs to improve developer productivity and streamline the software development process.
LLMs have achieved tremendous success in Code generation [52, 104]; however, there is a growing concern about the impact of software development on the environment. Training and deploying LLMs incurs significant environmental costs, including substantial CO2 emissions and water usage [34]. According to [86], training LLaMA 3.1 with 8 billion parameters generated approximately 420 tCO2e, which is equivalent to the emissions from 83 years of electricity usage by a single U.S. household. The process also consumed 2,769 kiloliters of water, roughly equal to 24.5 years of water usage by an average American, and this is only for the training phase. Once deployed, these models continue to consume energy as users interact with them. Energy usage during inference has increased rapidly [34], and total emissions depend on how frequently the model is used. For example, if ChatGPT receives 100 million queries per day [34], and each query consumes about 0.002 kWh of energy, the total daily energy consumption would be approximately 0.2 gigawatt-hours (GWh) [129].
Although LLM inference consumes a considerable amount of energy, the environmental impact does not end at the inference stage. Once integrated into software systems, the efficiency of the generated code itself becomes a critical factor in the application’s sustainability over its entire life-cycle [82]. Inefficient code can increase computational load, memory usage, and execution time, thereby raising energy consumption during operation. This is particularly relevant given that the information and communications technology (ICT) sector, which includes software and the supporting hardware infrastructure, currently accounts for an estimated 2–7% of global greenhouse gas emissions and is projected to rise to 14% by 2040 [42].
The growing integration of AI-assisted code generation tools further underscores the need to examine the sustainability of generated code. A recent large-scale study trained a neural classifier to detect AI-generated Python functions in 80 million GitHub commits (2018–2024) from 200,000 developers worldwide [30]. By December 2024, AI-generated code accounted for 30.1% of Python functions authored by U.S. contributors, 24.3% in Germany, 23.2% in France, 21.6% in India, 15.4% in Russia, and 11.7% in China, with newer GitHub users adopting AI tools more readily than experienced developers. Given this rapid adoption, it is increasingly important to investigate whether LLMs can produce not only correct and functional code, but also code that is sustainable throughout its operational life.
While a growing number of studies have examined the sustainability of LLM-generated code, the literature still lacks a detailed and systematic review on this topic. To address this gap, we present a systematic literature review (SLR) that identifies, collects, and organizes research on the sustainability of code generated by LLMs. Specifically, we review four types of contributions. First, we investigate studies evaluating how effectively LLMs produce code that aligns with sustainability principles in practice. Second, we examine the metrics and parameters used to assess sustainability, including measures of energy efficiency, resource usage, and carbon footprint. Third, we identify the benchmarks, datasets, and tools used to evaluate sustainability in LLM-generated code. Finally, we explore research analyzing how prompting strategies or fine-tuning methods influence the sustainability of code output.
To collect and review the relevant literature, we follow a systematic and structured process. We begin by identifying an initial set of studies through database searches and reference checking. We then apply clear inclusion and exclusion criteria in multiple screening stages to select the most relevant papers. For the selected studies, we extract data using predefined forms to ensure consistency and accuracy. We organize the collected information into tables and visual summaries to compare key characteristics of the reviewed studies. Based on this analysis, we identify common patterns, main findings, and existing limitations in the current body of research.
Overall, we find that research in this domain remains limited. The diversity of evaluated models is narrow, with most studies focusing on large language models, while small language models receive much less attention. Energy measurements are primarily conducted at the software level, and only a few studies use external hardware tools for a more accurate assessment. In addition, most research focuses on a small set of programming languages and general-purpose software development, with limited attention given to other software domains. We also observe the absence of dedicated sustainability-focused benchmarks designed specifically to evaluate the energy efficiency of LLM-generated code. Finally, sustainability-aware fine-tuning remains largely unexplored and represents an important direction for future research.
By synthesizing the current body of research, we provide a clear overview of the models studied, the programming languages considered, the evaluation methods applied, and the limitations that remain. This structured analysis helps identify research gaps and offers direction for future work toward more sustainable and environmentally responsible code generation using LLMs.
The remainder of this paper is organized as follows. Section II provides background information on LLMs and code generation. Section III describes the research methodology. Section IV provides insights into the extracted data, while Section V presents the analysis and results from the extracted data. Section VI presents the future research in the area. Section VII outlines the validity threats and limitations. Finally, Section VIII presents concluding remarks.
II Background
II-A Large Language Models
LLMs [87, 48] are advanced deep neural networks trained on large corpora of text (and, for code models, source code) to model the probability of token sequences. At inference time, an LLM takes previously seen tokens and predicts the next token. Given a sequence $x=(x_{1},\dots,x_{T})$, the model parameterized by $\theta$ factorizes the sequence likelihood as
$P(x)\;=\;\prod_{t=1}^{T}P(x_{t}\mid x_{<t};\theta).$
where $x_{t}$ is the token at position $t$, $x_{<t}=(x_{1},\dots,x_{t-1})$ are the preceding tokens, $\theta$ are the learnable model parameters.
General-purpose LLMs [122, 93] are trained primarily on large collections of text, with a smaller portion of code and mathematical content to enhance logical reasoning [68]. In contrast, code LLMs are pre-trained, or further trained, on large-scale code corpora, often with some text and math data, to specialize in programming-related tasks [68]. Some models, e.g., Qwen2.5-Coder[56], also make use of synthetic data to expand their training sets.
Architecturally, an LLM is a stack (concatenation) of Transformer blocks [131]. To pass the text data in the transformer, each input token (word) $x_{t}$ is first mapped to a continuous embedding [7] $e_{t}\;=\;E(x_{t})\in\mathbb{R}^{d}$ where $E$ is the embedding matrix and $d$ is the hidden size. A Transformer block takes the current hidden states and applies self-attention followed by a position-wise feed-forward network (with residual connections and layer normalization omitted here for brevity). This combination of self-attention and feed-forward layers, together with residual connections and normalization, forms the core building block of the Transformer architecture used in LLMs. For more information on transformer architectures and their underlying mechanisms, readers are referred to [131]
Transformer architectures [49, 94] are generally divided into three types: encoder-only, decoder-only, and encoder–decoder [20, 31]. Encoder-only models focus on understanding and analyzing input data. Decoder-only models are designed to generate sequences step by step, making them effective for creating new content. Encoder–decoder models combine both approaches, first processing the input and then generating the output, which is useful for tasks that involve transforming one type of data into another. Each architecture has its own strengths and is suited to different applications, depending on the nature of the data and the task [50].
II-A1 LLM Training
During training, LLMs learn to predict the next token in a sequence by optimizing their parameters $\theta$ over a large corpus of text or code. Given a sequence $x=(x_{1},\dots,x_{T})$, the model is trained using maximum likelihood estimation, which minimizes the negative log-likelihood [16]:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}\log P(x_{t}\mid x_{<t};\theta).$
The probability $P(x_{t}\mid x_{<t};\theta)$ is computed by first embedding each token into a continuous vector, processing the sequence through multiple stacked Transformer blocks, and finally projecting the hidden representation $h_{t}$ onto the vocabulary space via a linear transformation:
$z_{t}=W_{o}h_{t}+b_{o},\quad P(x_{t}\mid x_{<t};\theta)=\mathrm{softmax}(z_{t}),$
where $W_{o}$ and $b_{o}$ are learned output parameters. Gradients are computed via backpropagation through time (BPTT)[136], and the model parameters are updated using stochastic gradient descent or its variants (e.g., AdamW) [78].
II-A2 LLM Inference
During inference, the model generates tokens autoregressively. The user provides an initial prompt (context) $x_{1:k}$, which is tokenized and passed through the embedding layer and Transformer stack to obtain the hidden state $h_{k}$. This hidden state is used to compute the probability distribution over the vocabulary for the next token $x_{k+1}$:
$P(x_{k+1}\mid x_{1:k};\theta)=\mathrm{softmax}(W_{o}h_{k}+b_{o}).$
A decoding strategy such as greedy search (choosing the token with maximum probability), beam search [120], or sampling (e.g., top-$p$, top-$k$) selects the next token. This token is appended to the context, and the process is repeated. At each step, the model only attends to the tokens already generated (causal masking), ensuring that predictions are conditioned solely on past information:
$P(x_{k+1},\dots,x_{T}\mid x_{1:k})=\prod_{t=k+1}^{T}P(x_{t}\mid x_{<t};\theta).$
This iterative process continues until a termination condition is met, such as producing an end-of-sequence token or reaching a maximum length. In the context of code generation, the output tokens correspond to syntactically valid source code that fulfills the natural language or partial code prompt provided by the user.
II-A3 Small Language Models
Small Language Models (SLMs) share the same basic Transformer-based architecture as LLMs but have far fewer parameters, often meaning fewer transformer blocks or smaller layer sizes [80]. Their reduced size leads to lower computational and memory requirements during inference, making them faster and more suitable for deployment in resource-constrained environments [73]. SLMs can be trained from scratch on a specific dataset or created through model distillation [45], where a smaller model learns from a larger one. Although they may not match the performance of very large models on all tasks, well-trained SLMs can perform effectively in focused domains. In code-related tasks, they can provide competitive results in areas such as code completion, bug detection, and summarization, while offering faster inference and lower environmental impact[135].
II-A4 Energy Consumption
LLM inference operates in an auto-regressive manner, meaning the model generates one token at a time, with each new token depending on all the tokens that came before it, both from the original input prompt and from the tokens it has already generated [117]. First, the model processes the input prompt through its transformer layers to create a context. At each step, it appends the newly generated token to this context and reprocesses the entire sequence to predict the next token. An example of auto-regressive token generation is given below.
Autoregressive Token Generation Given the prompt “Write a Python function to add two numbers”, an LLM generates code in an autoregressive manner, producing one token at a time while reprocessing the entire sequence at each step: 1. Step 1: Input: “Write a Python function to add two numbers” Output token: def 2. Step 2: Input: “Write a Python function to add two numbers def” Output token: add_numbers 3. Process continues until the full function is generated. At each step, the model does not just consider the latest token; it reprocesses the entire prompt along with all previously generated tokens to produce the next token.
Because the model must process the full sequence at every step, the computational workload and memory usage grow with the total context length (input ++ output tokens) [134]. Longer prompts and longer outputs, therefore, require more energy. This effect is amplified in larger models, as each step involves executing billions of parameters through matrix multiplications, attention operations, and memory transfers[113]. Consequently, inference energy consumption scales with the model size, input length, and output length, with long code or text generations being significantly more costly than shorter ones. The energy is primarily consumed by GPUs performing these repeated, computationally intensive operations for each token generated [146].
In contrast, SLMs follow the same auto-regressive process but with far fewer parameters and a lighter architecture. This means each inference step requires fewer matrix multiplications, less attention computation, and reduced memory transfers. As a result, the energy consumed per generated token is significantly lower compared to large models. Additionally, smaller model sizes reduce the hardware requirements, allowing inference to be performed on less power-hungry devices, further improving energy efficiency for tasks such as code generation[90].
II-A5 Fine-tuning and Prompt Engineering
Fine-tuning [132] is a traditional method for adapting pre-trained language models to specific tasks by updating their internal parameters using task-specific datasets. During this process, the model is trained on labelled examples from the target domain, enabling it to learn patterns and vocabulary. While fine-tuning can achieve high performance, it requires significant computational resources, large amounts of labelled data, and can be time-consuming, particularly for large-scale models[118].
Prompt Engineering [140] has emerged as an effective alternative to model fine-tuning, particularly with large language models containing billions of parameters. Instead of modifying model weights, prompt engineering guides the model at inference time by providing carefully designed natural language instructions, optionally accompanied by examples. This approach enables task adaptation without additional training, making it computationally efficient and flexible. By structuring the input appropriately, prompt engineering can significantly influence the quality, correctness, and behaviour of the output, including in code generation tasks[107].
Prompting strategies vary based on the amount and structure of guidance provided. In zero-shot prompting, the model receives only a task description without examples. One-shot and few-shot prompting include one or several examples, respectively, to demonstrate the desired input–output relationship. More advanced techniques, such as Chain-of-Thought (CoT) and Program-of-Thought (PoT), encourage the model to generate intermediate reasoning steps or structured program-like outputs before producing the final result. These techniques are particularly useful for complex reasoning and code-related tasks. For a more comprehensive overview of prompt engineering techniques, readers are referred to existing surveys and dedicated studies in this area [81].
TABLE I: Research Questions
RQ Description
RQ1 How effectively do LLMs produce code that aligns with sustainability principles in practice?
RQ2 What metrics or parameters are used to evaluate sustainability?
RQ3 What benchmarks, datasets, and tools are utilized to assess the sustainability of LLM-generated code?
RQ4 How do prompting strategies or fine-tuning methods influence the sustainability of generated code?
II-B Code Generation
Code generation is the automatic production of executable source code from inputs such as natural language descriptions, design documents, or existing code, to reduce manual effort and improve development efficiency. Traditional approaches, often based on rule-based systems, templates, or domain-specific languages, provide partial automation for well-defined tasks but face key limitations [14, 141, 59]: they have limited contextual understanding, struggle to produce complete and logically consistent code, and lack adaptability to diverse software development needs. As modern development increasingly involves open-ended requirements, rapidly changing APIs, and complex architectures, these limitations restrict the effectiveness of traditional code generation in real-world applications [36].
Beyond code generation, LLMs are also applied to related programming tasks such as code completion, code optimization, and code refactoring. Code completion [19] refers to predicting and suggesting the next tokens, statements, or code blocks based on partially written programs, thereby assisting developers during interactive coding and improving productivity. Code optimization [62] focuses on improving existing code by enhancing performance aspects such as runtime efficiency, memory usage, or energy consumption while preserving its original functionality. In contrast, code refactoring [121] involves restructuring or reorganizing code to improve readability, maintainability, and overall structure without altering its external behaviour. Although these tasks serve different purposes, they all contribute to improving software quality and efficiency. From a sustainability perspective, these tasks are particularly important because they directly affect resource usage and the long-term environmental impact of software systems.
III Research Methodology
III-A Overview
As a first step, we searched for existing systematic literature reviews and surveys related to LLMs and code generation in digital libraries such as Compendex, Inspec, Scopus, and Google Scholar. We identified several recent reviews covering related areas, including general LLM-based code generation [32, 13, 17, 5, 67, 145, 23, 65], LLMs in software engineering [54, 53, 39], code completion [57, 58], automated program repair [148, 142, 147], parameter-efficient fine-tuning for large code models [2, 51], domain-specific and low-resource code generation [70, 46], LLMs in programming education [99, 21, 75, 3], developer productivity [84, 98], and quality and efficiency of LLM-generated code [103, 95, 63, 119]. In addition, systematic mapping studies have examined the use of LLMs for generating and reviewing code [4, 111, 1, 108].
However, despite the growing number of reviews in these related domains, we did not find any SLR that specifically focuses on the sustainability, energy efficiency, or environmental impact of LLM-generated code. To the best of our knowledge, this study is the first SLR that systematically analyzes sustainability-aware code generation using LLMs.
We then designed our own SLR following the guidelines proposed by Kitchenham [72]. In the first stage, we defined a review protocol, which included: (a) research questions, (b) a search strategy with a search string and relevant paper repositories, (c) inclusion and exclusion criteria, (d) an iterative study selection process, and (e) a data extraction strategy for classifying the primary studies. The process went through several cycles of execution, evaluation, and refinement until we finalized the protocol, which is described in the remainder of this section.
TABLE II: Inclusion and Exclusion Criteria
Inclusion Criteria
The study explicitly uses, evaluates, or fine-tunes an LLM for code-related tasks.
The study addresses sustainability aspects related to code generation.
The study includes empirical evaluation of generated code.
The full text is accessible.
Exclusion Criteria
The study does not involve an LLM or focuses only on traditional ML/DL/NLP approaches.
The study does not assess sustainability or energy-related aspects.
The study focuses on tasks unrelated to code generation (e.g., text summarization, image captioning, question answering).
The study discusses efficiency but does not perform energy-related experimental evaluation.
The study addresses sustainability in domains unrelated to code generation (e.g., green supply chain management).
The study is purely theoretical or conceptual without empirical validation.
The publication is a doctoral symposium paper, editorial, position paper, or poster.
The study is a survey, review, or tutorial without original experimental results.
The full text is not accessible or is not written in English.
III-B Research Questions
The following research questions (RQs) were defined to guide the systematic literature review in Table I
RQ1 is motivated by the need to evaluate how well LLMs can generate code that meets sustainability goals, such as reducing energy consumption, lowering computational overhead, and minimizing resource usage. It examines whether sustainability is an inherent property of LLM-generated code or the result of deliberate design choices. The scope includes code generated across different application domains and programming languages. Addressing RQ1 will help researchers identify which LLMs produce more sustainable code and understand the factors that contribute to this outcome.
RQ2 aims to identify and categorize the measures used to evaluate the sustainability of LLM-generated code. These may include metrics such as energy consumption, carbon footprint, memory usage, execution time, and other efficiency indicators. Examining these metrics will show how sustainability is defined and assessed in existing studies. The findings will also support the development of standardized evaluation methods, enabling more consistent comparisons across future research.
RQ3 focuses on identifying the resources used to assess the sustainability of LLM-generated code, including benchmarks, datasets, simulation frameworks, and performance analysis tools. It examines whether these resources are specifically designed for sustainability evaluation or adapted from general-purpose code generation benchmarks. The findings will help researchers choose appropriate resources for experimentation and highlight gaps where new sustainability-oriented datasets or tools are needed.
RQ4 examines how model adaptation techniques affect the sustainability of generated code, including prompt engineering, instruction tuning, and domain-specific fine-tuning, etc. It focuses on the direct impacts of these techniques on LLM-generated code, such as reduced runtime and lower energy consumption. The findings will help establish best practices for designing prompts and fine-tuning methods that enhance sustainability without reducing functional correctness.
III-C Search Strategy
Guided by the RQs, we adopt an iterative search strategy. We repeatedly query digital libraries using Boolean combinations of relevant keywords, and after each iteration, we refine the keywords to improve coverage. This process continues until the final search results include all publications from a manually identified benchmark set.
Figure 1: Primary Study Selection
III-C1 Search String
Our search term was structured around two main aspects of the topic. The first aspect relates to LLMs, code generation, and associated techniques such as fine-tuning. The second aspect focuses on sustainability and energy efficiency. Combining these aspects allowed us to capture the broadest possible set of studies relevant to sustainable code generation using LLMs. The final search term was formulated by linking keywords from both aspects to ensure that retrieved publications addressed concepts from each domain.
We refined the search string iteratively by evaluating the results it produced. To guide this process, we manually compiled a benchmark set of thirteen (13) relevant research publications from multiple sources, including Google Search, Google Scholar, and arXiv. In the first iteration, a comparison between the benchmark set and the search results showed that only 10 out of the 13 publications were retrieved. To capture the three missing publications, we added additional terms such as “LLM-generated code,” “energy efficiency,” and “efficient code” to the search string. After these additions, the revised search string successfully returned a result set containing the entire benchmark set. This final search string was adopted for the remainder of the study and is presented in Table III.
III-C2 Search Sources
We searched for relevant publications in five databases: IEEE Xplore, Compendex, Inspec, Scopus, and the ACM Digital Library. The retrieved publications were managed using an Excel spreadsheet. A detailed breakdown of the search process for each database, along with the corresponding number of publications, is presented in Fig. 1.
TABLE III: Search String
(”large language model*” OR ”LLM” OR ”Small Language Model*” OR ”SLM” OR ”GPT” OR ”Codex” OR ”ChatGPT” OR ”code generation” OR ”code generation model*” OR ”AI code assistant” OR ”AI-assisted programming” OR ”automated programming” OR ”code refactoring” OR ”code optimization” OR ”code synthesis” OR ”program synthesis” OR ”code completion” OR ”software language model” OR ”AI programming tool*” OR ”generative AI for code” OR ”code LLM” OR ”AI-based code generation” OR ”prompt engineering” OR ”instruction tuning” OR ”fine-tuning” OR ”zero-shot learning” OR ”few-shot prompting” OR ”chain of thought” OR ”model adaptation” OR ”Llm-generated code”) AND (”sustainable code” OR ”green software” OR ”green software engineering” OR ”energy efficient” OR ”eco-friendly software” OR ”low-energy code” OR ”energy-aware” OR ”carbon footprint” OR ”environmental impact” OR ”code sustainability” OR ”energy-optimized software” OR ”energy-efficient software” OR ”sustainability in software engineering” OR ”sustainable programming” OR ”energy efficiency” OR ”efficient code”)
TABLE IV: Data Extraction Form
Category Data Items Description
General Context of Study Categorized as an academic or industrial setting.
Code-Related Task The primary coding task addressed (e.g., code generation, completion, etc.
Coding Application The application domain of generated code (e.g., blockchain, scientific computing, etc).
Model Evaluated Models and Size Categories Lists and classify the LLM(s) used in the study, along with their parameter sizes.
Model Openness Specifies whether the models are open or closed-source, or a combination of both.
Model Domain Specialization Indicates whether the model is general-purpose, code-specific, or a combination of both.
Evaluation Evaluation Metrics Metrics used to assess performance and sustainability.
Custom Evaluation Metric Indicates whether the study introduces a novel evaluation metric.
Libraries / Tools for Measurement Tools or libraries used for metric collection.
Hardware Configuration Description of the hardware and experimental setup used for code execution.
Experimental Configuration Details regarding experiment repetitions and inference parameters (e.g., temperature, top-p, maximum tokens, batch size, precision).
Open-Source Availability Availability of code, prompts, datasets, or other resources to support reproducibility.
Benchmark Usage Indicates whether established benchmarks are used for model evaluation.
Model Adaptation Fine-Tuning Details If the study is performing sustainability-aware fine-tuning.
Use of Prompt Engineering Specifies whether prompt engineering strategies are applied to generate sustainable code.
III-D Inclusion and Exclusion Criteria
To ensure that only the most relevant and high-quality studies were considered in this systematic literature review, we applied a predefined set of inclusion and exclusion criteria, as summarized in Table II. In order to qualify as a primary study, the research work must meet all inclusion criteria and must not meet any exclusion criteria.
We included studies that used or evaluated LLMs for code generation, addressed sustainability aspects such as energy efficiency or resource usage, and provided empirical evaluation results. Studies were excluded if they did not involve an LLM, lacked sustainability analysis, focused on unrelated tasks, or had no accessible full text in English. The inclusion and exclusion criteria were applied at multiple stages of the literature review process, including the initial filtering of primary studies and the full-text screening phase. They were also employed during the snowballing process, which resulted in the addition of 4 more primary studies.
Although several studies claim efficiency improvements, many of them do not provide empirical evaluation of energy consumption or energy efficiency. Since our objective is to examine the sustainability of LLM-generated code based on measurable evidence, we refined our exclusion criteria to remove studies that do not include any form of energy consumption analysis. As a result, only studies that provide explicit energy-related evaluation were considered in the final selection. This refinement reduced the number of eligible studies to 19. Moreover, to ensure comprehensive coverage of this emerging research area, we also included relevant pre-prints from arXiv, as sustainability-focused research on LLM-generated code is still developing and may not yet be fully represented in peer-reviewed venues.
Study Models Evaluated Model Size Category
[96] GPT-4o Large
[110] Llama 3 8B-Instruct Medium
[127] ChatGPT-3.5, ChatGPT-4o, CodeGemma-7B, WizardCoder-33B, Llama 3-8B, Phind-CodeLlama-34B-v2, Nous-Hermes 2-10.7B, Mistral-7B Mixed (Small + Large)
[33] Not Specified Not specified
[22] Code Llama-70B, Code Llama-70B-Instruct, Code Llama-70B-Python, DeepSeek-Coder-33B-Base, DeepSeek-Coder-33B-Instruct Large
[27] GitHub Copilot (StarCoder family back-end models: StarCoder 15.5 B, StarCoder2-7 B, StarCoder2-15 B) Mixed (Small + Medium)
[60] Llama 3.2 (3 B) and OPT (2.7 B) Small
[25] GPT-4 Large
[…]
[130] GitHub Copilot Not specified
[129] GitHub Copilot, OpenAI ChatGPT-3, Amazon CodeWhisperer Large
[64] Amazon Nova-Lite, Amazon Nova-Micro, Amazon Nova-Pro, Anthropic Claude 3.5 Haiku, Anthropic Claude 3.5 Sonnet, DeepSeek v3 (37B), Google Gemini 1.5 Flash, Google Gemini 1.5 Pro, Google Gemini 2.0 Flash, Google Gemini 2.0 Flash-Lite, Meta Llama 3.1 (8B), Meta Llama 3.1 (70B), Meta Llama 3.3 (70B), Mistral Codestral-Mamba-2407 (7B), Mistral-Large-2407 (123B), Pixtral-Large-2411 (124B), OpenAI GPT-3.5 Turbo, OpenAI GPT-4 Turbo, OpenAI GPT-4o, xAI Grok Mixed (Small + Medium + Large)
[101] GitHub Copilot Not specified
[11] StableCode-3B, StarCoderBase-3B, Qwen2.5-Coder-3B-Instruct, GPT-4.0, DeepSeek-Reasoner Mixed (Small + Large)
[97] GPT-4o, GPT-4.1, qwen3-coder:480b, gemma3:27b, deepseek-r1:70b, llama4:latest Mixed (Medium + Large)
[116] GitHub Copilot Not specified
[29] CodeLlama-7B-Instruct Small
[12] StableCode-Instruct-3B, Qwen2.5-Coder-3B-Instruct, CodeLlama-7B-Instruct, Phi-3-Mini-4K-Instruct Small
[10] ChatGPT (175B), GPT-4 (1.76T), DeepSeek Coder Instruct (33B), Speechless CodeLlama v2.0 (34B), CodeMillennials (34B), WizardCoder v1.1 (33B) Large
TABLE V: Overview of the evaluated models and their parameter-scale categories.
III-E Study Selection Procedure
After applying the search string (Table III), we initially retrieved 4,327 publications. Removing duplicates reduced this number to 2,015. This refined set was then screened by title and abstract to identify the initial pool of primary studies. To improve the reliability of this manual filtering process, we iteratively refined the inclusion and exclusion criteria. First, we randomly selected a sample of 100 publications from the 2,015 retrieved studies. Each author independently reviewed these papers against the criteria and decided whether they should be included in the initial pool. We then measured inter-rater agreement using Fleiss’ Kappa [40]. If the agreement was low, the authors discussed the disagreements, revised the criteria, and repeated the process until an acceptable level of inter-rater reliability was achieved.
Three such rounds emerged. Our initial Fleiss’ Kappa index was 0.378, indicating low agreement. Post-analysis showed that the main issues were unclear selection criteria and inconsistent interpretations of sustainability-related concepts. Disagreements arose from including papers on energy efficiency in domains other than code generation, as well as studies on DNNs or CNNs, which were outside our scope. Another challenge was the definition of sustainability, where some studies focused solely on efficiency without explicitly addressing sustainability. We refined the inclusion criteria to focus strictly on LLMs and code generation by excluding studies on general neural networks and other application domains, and then conducted a second iteration. In the second iteration, Fleiss’ Kappa inter-rater agreement increased to 0.833, indicating that the researchers had reached a sufficient and consistent understanding of the inclusion criteria.
We then applied the refined inclusion criteria independently to the entire set of retrieved publications, identifying 170 studies for full-text review. To maintain consistency, we used the same pilot process with Fleiss’ Kappa measurement during this stage. After carefully reading the full texts, we found that only 15 studies fully met the refined criteria and were included as primary studies. We chose not to apply additional quality assessment filters, as the final pool of studies was already small and the research area is still emerging. This approach allowed us to remain as inclusive as possible while ensuring relevance to our research focus.
We conducted a snowballing [66] search to identify additional relevant primary studies. Snowballing involves reviewing the reference lists of papers (backward snowballing) and the papers that cite them (forward snowballing) to find further studies. In this study, both forward and backward snowballing were applied to the 15 retrieved primary studies. Google Scholar was used to obtain the citations for each paper, and the same selection criteria were applied to both their references and citations. This process yielded 4 additional relevant papers, bringing the final pool of primary studies to 19. Figure 1 shows how the initial 2,015 publications retrieved from the databases were reduced to these 19 primary studies, which were then used for data extraction.
III-F Data Extraction Form
We created a structured data extraction form to collect information from each primary study consistently and accurately, focusing on our research questions (RQ1–RQ4). First, three researchers prepared a draft extraction form based on a small sample of studies. We then met to combine these drafts into the final extraction form shown in Table IV. The categories in the form were designed to capture the main differences found in research on sustainable code generation using LLMs.
The General category records basic information such as the study setting, the main code-related task, and the application area. The Model category records details about the LLMs used, including whether they are open or closed source, their size, domain specialization, and other identifying features. The Dataset / Benchmark category records whether a study used a standard coding benchmark or a custom dataset, along with details such as dataset source, programming languages, number of problems, and baselines used. The Evaluation category records the metrics, tools, hardware, and configurations used to measure performance and sustainability, as well as reproducibility information. The Model Adaptation category records whether and how the base LLM was adapted before or during evaluation, including fine-tuning and prompt engineering methods, and any reported effect on sustainability metrics.
The extracted data offers a clear framework for analysis, enabling the comparison and summarization of findings across all studies. The results are presented in the next section.
IV Data Extraction
We present the analysis of data collected from the final pool of 19 primary studies. This section outlines the extracted data, as summarized in Table IV, providing a structured basis for analysis and enabling comparisons across all studies. The information is then used in Section LABEL:subsec:results_rqs to synthesize findings and address RQ1–RQ4.
IV-A General
IV-A1 Context of Study
We first categorized the selected studies based on the environment in which the research was undertaken, distinguishing between academic and industrial settings. A study was labelled as academic if all contributing authors were affiliated with universities or research institutions. A study was labelled as industrial if the research involved participation from industry, identified by the presence of at least one author with an industrial affiliation. The distribution of studies across these contexts is summarized in Table VI. Out of the 19 studies, 14 were conducted in an academic setting, while only 5 were performed in an industrial setting.
Infrastructure Type Primary Studies Total
Academic [96], [110], [127], [33], [22], [27], [60], [25], [105], [101], [97], [116], [29], [10] 14
Industrial [130], [129], [64], [11], [12] 5
TABLE VI: Context of study
IV-A2 Code-Related Task
We categorized the selected studies based on the code-related tasks they addressed. Code-related tasks refer to code generation, code completion, code optimization, and code refactoring. As several studies addressed more than one task (e.g., combining code generation with optimization or refactoring), a multi-label classification approach was adopted; therefore, a single study may appear in multiple task categories. This approach enables a more accurate representation of the scope of each study and avoids oversimplifying contributions that span multiple tasks.
Infrastructure Type Primary Studies
Code Generation [33], [127], [22],[27], [60], [25], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10]
Code Optimization [96], [105], [101], [116]
Code Completion [110]
Code Refactoring [33]
TABLE VII: Code-related tasks
Table VII shows the distribution of studies across code-related tasks. It can be seen that the majority of the selected studies focus on code generation, indicating that energy-aware research in this area is primarily centred on generating source code using LLMs. In contrast, relatively fewer studies address code optimization, code completion, or code refactoring, suggesting that these tasks have received limited attention from an energy measurement perspective.
IV-A3 Code Application Domain
To better understand the application contexts of the selected studies, we categorized them based on the type of software and workloads they target. General-purpose software development includes programming tasks implemented using standard programming languages and execution environments, such as algorithmic problem solving and common application development, which are not tied to a specific scientific or engineering domain. In contrast, scientific and engineering applications focus on domain-specific workloads involving numerical computation, simulation, or parallel processing. The distribution of studies across these application contexts is presented in Table VIII.
Infrastructure Type Primary Studies Total
General-Purpose Software Development [96], [110], [127], [22], [27], [60], [25], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10] 17
Scientific and Engineering Applications [105], [33] 2
TABLE VIII: Application domains targeted by the selected studies
It can be observed that only two studies focus on scientific and engineering applications. In [33], the authors propose LASSI-EE, an automated LLM-based refactoring framework that generates energy-efficient CUDA code [114]. In [105], the authors evaluate the effectiveness of LLMs in reducing the environmental footprint of real-world MATLAB projects.
IV-B Model
IV-B1 Evaluated Models and Model Size Categories
We first present the list of models considered in each study in Table V. The models are listed as reported by the original authors to accurately reflect the experimental settings used in each work. One study does not explicitly report the underlying models employed and is therefore marked as Not Specified.
To enable consistent comparison across studies, we also categorize the evaluated language models based on their parameter count into three size categories: small, medium, and large. Small models have at most 7 billion parameters ($\leq 7B$), medium models have between 8 and 30 billion parameters, and large models exceed 30 billion parameters ($\geq 30B$). Some studies evaluate models spanning multiple size categories and are therefore labelled as mixed-scale to accurately reflect their evaluation scope. Studies that do not report model sizes or rely on proprietary systems with undisclosed back-end configurations are labelled as not specified. This classification enables a clear and reproducible analysis of model usage trends while accounting for differences in computational scale.
IV-B2 Model Openness
To further characterize the experimental settings of the selected studies, we categorized the evaluated models based on their availability and licensing. Studies were classified as closed-source if they exclusively evaluated proprietary or commercial models, open-source if they evaluated only publicly available models, and both open-source and closed-source if they considered a combination of proprietary and open-source models.
Category Primary Studies Total
Open-Source [110], [22], [27], [60], [29], [12] 6
Closed-Source [96], [25], [130], [129], [101], [116] 6
Both Open and Closed-source [127], [105], [64], [11], [97], [10] 6
Not Specified [33] 1
TABLE IX: Categorization of studies based on model availability
Table IX summarizes the categorization of the selected studies based on model openness. It can be seen that the studies are evenly distributed across open-source, closed-source, and mixed model categories, with six studies in each group. In addition, although one study [33] discusses the use of proprietary models, it does not clearly specify which models are evaluated and is therefore categorized as not specified.
IV-B3 Model Domain Specialization
To characterize the domain focus of the evaluated models, we classify them as general-purpose, code-specialized, or mixed (general and code-specialized) language models. General-purpose LLMs [133] are designed for a broad range of natural language tasks without domain-specific optimization. Code-specialized LLMs [9] are trained or fine-tuned for programming-related tasks, including code generation, completion, and analysis. Studies that evaluate both general-purpose and code-specialized models are classified as mixed. Studies that do not report the domain specialization of the evaluated models are labelled as not specified.
Category Primary Studies Total
General-Purpose LLMs [96], [110], [60], [25], [105] 5
Code-Specialized LLMs [22], [27], [130], [101], [116], [29], [12] 7
Mixed [127], [129], [64], [11], [97], [10] 6
Not Specified [33] 1
TABLE X: Categorization of studies based on model domain specialization
The distribution of studies across these domains is presented in Table X. It can be seen that only five studies focus exclusively on general-purpose language models, while the remaining studies evaluate either code-specialized models or include at least one code-specialized model in a mixed evaluation setting.
IV-C Evaluation
IV-C1 Evaluation Metrics
We categorize the evaluation metrics considered in the primary studies into six groups: energy, performance, correctness, memory, code quality, and miscellaneous metrics. Energy metrics capture absolute and relative energy usage as well as environmental impact. Performance metrics [137] measure execution time, latency, and throughput. Correctness metrics assess the functional validity of generated outputs. Memory metrics [88] quantify resource utilization, including memory usage and computational operations. Code quality metrics [115] evaluate lexical, structural, and semantic similarity between generated and reference code. Miscellaneous metrics capture operational and cost-related aspects that are not covered by the other categories. For each metric, the table lists the primary studies in which it is used, providing a structured overview of evaluation practices in the literature.
Table XI summarizes the evaluation metrics reported across the selected studies. It can be observed that most studies focus on energy consumption and performance metrics, particularly latency and runtime. Fewer studies evaluate functional correctness and memory usage, and only a small number consider code quality.
Category Evaluation Metric Primary Studies
Energy Energy consumption [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10]
Power consumption [27], [12]
EnergyRatio [127]
Energy-reduction@k [33]
Carbon footprint [25]
Performance Latency/Runtime [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [12]
Runtime Ratio (%) [127]
Throughput (tokens/s) [60], [97]
CPU cycles [97]
CPU usage (%) [116]
Correctness Functional correctness [96], [105], [130], [129], [11], [97]
Pass@k [33], [64]
Memory Memory usage [105], [130], [129], [64], [11], [97], [116], [12]
Peak memory [22]
FLOPs [22], [130], [129]
Code Quality Exact Match [110]
CodeBLEU, ROUGE-L [60]
Syntax & DataFlow score [60]
AST similarity (%) [22]
Miscellaneous Token usage [64]
Monetary cost [64]
Rejected requests [27]
Completed generations [27]
TABLE XI: Categorization of evaluation metrics reported in the selected studies.
IV-C2 Custom Evaluation Metrics
In addition to standard evaluation metrics, a small number of studies introduce custom sustainability-oriented metrics. These metrics are designed by the authors to capture specific aspects of energy efficiency, carbon impact, or sustainability that are not covered by conventional measures. Table XII presents the studies that propose such custom evaluation metrics and lists the corresponding metrics introduced in each work. Studies that do not propose any custom evaluation metrics are omitted from this table.
Study Custom Evaluation Metrics
[127] RuntimeRatio, EnergyRatio
[33] Energy-reduction@k
[130] Embodied Energy, Operational Energy, Carbon Intensity, Embodied Carbon Emissions, Operational Carbon Emissions
[129] Green Capacity (GC)
TABLE XII: Studies proposing custom sustainability-oriented evaluation metrics.
RuntimeRatio and EnergyRatio compare the runtime and energy consumption of model-generated code against canonical human-written solutions, providing a relative efficiency measure. Energy-reduction@k estimates the expected energy savings when generating multiple candidate solutions and selecting the most energy-efficient valid one, while accounting for both performance improvement and correctness probability. Similarly, Green Capacity (GC) measures the overall sustainability improvement of optimized code across multiple performance dimensions, filtering out invalid or degraded solutions. Embodied Energy and Embodied Carbon capture the environmental impact associated with code generation or development infrastructure, whereas Operational Energy and Operational Carbon measure emissions during actual code execution. Carbon Intensity (CI) quantifies the amount of carbon emissions per unit of electricity consumed.
IV-C3 Libraries / Tools for Measurement
To understand how energy is measured in the selected studies, Table XIII presents the tools and libraries used for measuring energy consumption and related system metrics. The table lists each tool as reported by the original authors and maps it to the corresponding primary studies. It can be observed that most studies rely on Python-based measurement libraries or operating system–level tools, such as perf and Intel RAPL. In contrast, only two studies [29, 10] employ external hardware power monitors (Monsoon Power Monitor), indicating limited use of direct physical energy measurement in the existing literature.
Tool / Library Primary Studies
Intel RAPL[61] [96], [97]
CodeCarbon[28] [110], [25], [11], [12]
PyJoules[15] [127]
Perf[100] [22], [27], [129], [64], [116]
PyNVML[91] [33], [60]
Nvidia-smi[92] [27]
Rocm-smi[109] [33]
Turbostat[126] [101]
EnergiBridge[112] [105], [10]
Windows E3[83] [130]
Monsoon Power Monitor[85] [29], [10]
TABLE XIII: Tools and libraries used for energy measurement
IV-C4 Hardware Configuration
Table XIV summarizes the hardware environments used for evaluation in the selected studies. To improve clarity and comparability, individual hardware specifications are grouped into high-level evaluation categories based on where and how experiments are executed. Each category lists the primary studies that employ the corresponding evaluation setup. On-premise (local machine/server) refers to evaluations conducted on locally owned or institution-managed machines and servers where hardware resources are directly controlled by the authors. Cloud VM [38] / Cloud bare metal [149] includes evaluations performed on cloud-based virtual machines or dedicated cloud servers provided by commercial or academic cloud platforms. Containerized environment denotes evaluations executed within container-based setups; for example, the authors in [105] run their experiments using Docker containers [35]. Edge / embedded device [74] corresponds to evaluations carried out on low-power or embedded platforms; for instance, the authors in [29, 10] conduct their experiments on Raspberry Pi devices [128]. Finally, not specified indicates studies that do not explicitly report the hardware used for evaluation, which limits reproducibility and comparability.
Evaluation Hardware Category Representative Configurations
On-premise (local machine/server) [110], [127], [22], [27], [60], [129], [64], [116], [10]
Cloud [101], [11], [97], [12]
Containerized Environment [105]
Embedded Device [29], [10]
Not Specified [96], [33], [25], [130]
TABLE XIV: Categorization of hardware configurations used for evaluation in the selected studies.
IV-C5 Experimental Configuration
Table XV summarizes the experimental configuration, including the number of runs, decoding strategy [139], temperature [125], maximum token limits, and top-p [89] and top-k [143] values. The table reports these parameters exactly as described by the original authors to reflect their experimental setups.
Only studies that explicitly report at least one experimental or inference parameter are included in the table. Several studies partially specify their configuration by reporting parameters such as the number of runs or temperature, while leaving other settings unspecified. Overall, the table highlights substantial variation in experimental configurations and reveals that many studies provide limited details on inference settings, which may affect reproducibility and comparability across results.
Study Runs Decoding Strategy Temperature Max Tokens Top-p
[110] 10 - - - -
[127] 5 - 0.7 1024 -
[33] 30 - 0.2 - -
[22] 50 Greedy 0.2 - -
[27] 5 - - - -
[60] - - 0.6-0.92 15 -
[25] 3 - - - -
[105] 30 - 0.5 4096–8192 1
[129] 10 - - - -
[64] 25 - - - -
[11] 10 - - - -
[97] 5 - 0.7 - -
[116] 10 - - - -
[29] 30 - 0.6-1 - -
[12] 10 - - - -
[10] 30 - - - -
TABLE XV: Decoding and inference settings reported in the selected studies.
IV-C6 Open-Source Availability
To assess reproducibility and transparency, we examine the open-source availability of the selected studies. A study is considered open source if it provides publicly accessible code or a replication package. Table XVI summarizes the open-source availability of the reviewed studies, indicating whether implementation artifacts are publicly released.
Category Count
Open-source available [96], [110], [33], [22], [27], [60], [25], [105], [64], [97], [116], [29], [10]
Not open-source [127], [130], [129], [101], [11], [12]
TABLE XVI: Summary of open-source availability
IV-C7 Benchmark Comparison and Usage
We also categorize the evaluation data based on the type of benchmark or dataset used and the programming language considered. Table XVII summarizes the benchmarks and datasets used in the selected studies, along with programming language information when explicitly reported. Programming languages are reported only when explicitly specified by the authors or when the language choice is central to the evaluation; for standardized benchmarks, the programming language is implicit and therefore omitted. It can be observed that most studies primarily focus on Python, reflecting its widespread use in code generation research. In addition, not all studies rely on standardized benchmarks, as several works evaluate models using diverse or custom data sources, indicating variability in evaluation practices across the literature.
Study Benchmark / Dataset Source Programming Language
[96] Energy-Language[37] C++
[110] CodeXGLUE[79]
[127] LeetCode[76] Python
[33] HeCBench suite[69], XSBench[41], miniMDock[123]
[22] LeetCode Python
[27] CLI Connect 4 Game[124] Java
[60] JavaCorpus[6], PY150[106] Java, Python
[25] Codeforces[26] Python
[105] GitHub repositories[44] MATLAB
[130] Not specified
[129] LeetCode Python
[64] EffiBench[55]
[101] Not Specified Java
[11] LeetCode Python
[97] HumanEval_CPP[151], SciMark2[102], DaCapoBench[18]
[116] HumanEval[24]
[29] LeetCode / HackerRank[47] C++, JavaScript, Python
[12] LeetCode Python
[10] EvoEval [138]
TABLE XVII: Benchmarks and programming languages used in the selected studies.
IV-D Model Adaptation
IV-D1 Fine-Tuning
We also categorize the studies based on the use of fine-tuning, that is, whether a study adapts the model parameters to generate more sustainable code. Surprisingly, only one study [60] employs fine-tuning for this purpose. This study proposes a fine-tuning approach for code generation models using a weighted aggregated loss [77], which enables models to perform dynamic early exits at different layers without requiring additional language model heads. In addition, it introduces GREEN-CODE, a reinforcement learning–based [71] framework that improves energy efficiency in code generation by dynamically selecting early exit points during inference.
IV-D2 Prompt Engineering
Table XVIII summarizes whether the selected studies use prompt engineering techniques. A study is marked as “Yes” if it explicitly applies methods such as zero-shot, few-shot, chain-of-thought, iterative prompting, or prompt variations to influence code generation. Studies that use default model settings without modifying prompts are marked as “No.” It can be observed that several studies rely on prompt-based strategies, while others evaluate models without applying specific prompt engineering techniques.
Prompt Engineering Primary Studies
Yes [96], [110], [33], [22], [27], [101], [97], [29], [12], [10]
No [127], [60], [25], [105], [130], [129], [64], [11], [116]
TABLE XVIII: Use of prompt engineering techniques in the selected studies.
V Future Research Opportunities
We now present general observations and insights from our systematic review on sustainable code generation using LLMs, highlighting key research gaps and directions for future work.
V-A Expanding Sustainability Beyond Code Generation
One notable observation is the need to extend sustainability evaluation beyond code generation tasks. As shown in Table. VII, most existing studies focus primarily on generating new source code using LLMs, while relatively few examine code optimization, code completion, or code refactoring. This imbalance suggests that energy-aware research remains narrowly concentrated on initial code creation.
Future research should pay greater attention to optimization, completion, and refactoring tasks, as these activities play a critical role in improving software efficiency and long-term sustainability. Code optimization directly affects runtime performance, memory usage, and energy consumption, making it essential for reducing computational waste. Code refactoring improves structural quality and maintainability, which can indirectly enhance energy efficiency by eliminating redundant or inefficient logic. Similarly, intelligent code completion can guide developers toward more efficient implementations early in the development process, preventing inefficient patterns before deployment. Since software sustainability depends not only on generating code but also on improving and maintaining it over time, these areas represent significant and currently underexplored opportunities for advancing energy-aware LLM research.
V-B Underexplored Application Domains
An important gap identified in this review is the limited exploration of sustainability-aware code generation across diverse application domains. Majority of the selected studies focus on general-purpose software development, with only two studies addressing scientific and engineering applications. This distribution indicates that sustainability-aware research in LLM-based code generation is primarily centered on standard programming tasks and general development scenarios.
Future research should explore sustainability-aware code generation across a broader range of application domains. In particular, mobile and client-side applications, low-level and systems software, web and cloud infrastructure, and IoT and embedded systems are not adequately represented in the selected studies. These domains are especially important from a sustainability perspective, as energy efficiency directly affects battery life, system performance, scalability, and overall environmental impact. For example, embedded and IoT systems operate under strict energy constraints, low-level software can significantly influence hardware efficiency, and cloud-based services contribute substantially to large-scale energy consumption. The limited attention to these areas highlights the need for future studies to evaluate LLM-generated code in more realistic, energy-sensitive, and deployment-oriented environments.
V-C Expanding Sustainability Evaluation Across Programming Languages
Another important direction for future research is the exploration of sustainability-aware code generation across a broader range of programming languages. Most of the selected studies focus primarily on high-level languages such as Python, Java, and C++, with only one study examining MATLAB. Although these languages are widely used in both research and industry, they do not fully capture the diversity of real-world software systems and deployment environments.
This direction is particularly important from a sustainability perspective, as energy efficiency can vary significantly across programming languages and execution environments. Different languages introduce different abstractions, memory management mechanisms, and runtime behaviors, all of which influence energy consumption. Moreover, language choice is often tightly coupled with the target platform, such as mobile devices, cloud infrastructure, edge systems, blockchain networks, or networked distributed systems, which further shapes optimization opportunities, runtime characteristics, and energy profiles.
Low-level and system-oriented languages such as C and Rust provide fine-grained control over memory management, concurrency, and hardware interaction, enabling highly optimized and potentially energy-efficient implementations, although they require careful design and expertise. Evaluating LLM-generated code in such languages would offer deeper insight into whether models can produce efficient system-level solutions. Similarly, languages such as Solidity and Rust are widely used in blockchain development, where computational efficiency has direct financial and environmental implications. Smart contracts written in Solidity execute across distributed blockchain networks, and inefficient implementations can increase gas consumption and contribute to higher energy usage at scale. Beyond these examples, future research should also examine other programming environments, including Go for cloud-native systems, JavaScript for client-side applications, Swift and Kotlin for mobile development, and domain-specific languages used in scientific computing and data engineering. Expanding sustainability evaluation across diverse programming ecosystems would provide a more comprehensive and realistic understanding of energy-aware code generation in real-world deployment scenarios.
V-D Toward Sustainability-Aware LLM Optimization
Another important direction for future research is the development of sustainability-aware fine-tuning and model adaptation techniques. Our review shows that only one study [60] applies fine-tuning or reinforcement learning–based adaptation with an explicit focus on improving energy efficiency. This suggests that most existing work measures sustainability at the output level without modifying the underlying model to generate more energy-efficient code.
Future research should explore energy-aware fine-tuning strategies specifically tailored for sustainable code generation and optimization tasks. Instead of only measuring the energy consumption of generated code after inference, models should be trained to directly produce energy-efficient implementations. For example, training objectives can be extended to penalize inefficient code patterns, high-complexity logic, or unnecessary computational operations, encouraging the generation of optimized and resource-efficient solutions. Green reinforcement learning approaches may also be applied, where models receive reward signals based on improvements in runtime efficiency, reduced energy consumption, or lower carbon footprint of the generated code. Moreover, parameter-efficient adaptation methods can further reduce training and deployment cost, making sustainability-aware code generation more practical. By embedding sustainability objectives directly into model training and inference processes, future research can enable LLMs to proactively generate optimized, energy-efficient code rather than relying solely on post-hoc energy evaluation.
V-E Limited Focus on Small Models
Another important observation from our review is the limited focus on small language models in sustainability-aware code generation research. Only three studies exclusively evaluate small models, while four additional studies include small models alongside medium and large models for comparative analysis. This indicates that the majority of existing research remains centered on large-scale models, with relatively limited dedicated investigation of lightweight architectures.
Small language models generally require less computational power and consume less energy during inference, making them promising candidates for sustainable and deployment-oriented code generation, especially in edge and resource-constrained environments. However, there is still limited understanding of how these models perform with respect to correctness, efficiency, and the trade-off between energy consumption and accuracy under controlled experimental conditions. Most existing studies treat small models as secondary baselines rather than as primary subjects of investigation.
Future research should conduct systematic and controlled evaluations of small models to better assess their practical suitability for energy-efficient code generation. In addition, studies may explore techniques such as parameter-efficient adaptation, lightweight fine-tuning, quantization, or prompt optimization to improve performance while maintaining low energy consumption. A deeper understanding of the accuracy–energy trade-off is essential for designing sustainable AI systems that achieve an effective balance between performance, resource utilization, and environmental impact.
V-F Need for Sustainability-Oriented Benchmarks
Another important observation from this study is that none of the selected works introduce a benchmark specifically designed to evaluate sustainability. Instead, all studies use existing general-purpose benchmarks and add sustainability-related metrics, such as energy consumption or carbon footprint, as additional evaluation measures. This indicates a clear research gap in the field. Future research should focus on developing standardized benchmarks dedicated to sustainability-aware code generation. These benchmarks should include tasks and evaluation methods that measure energy efficiency, carbon impact, resource usage, and the ability of models to produce energy-efficient solutions. Creating such benchmarks would improve comparability across studies and support more consistent and rigorous evaluation of sustainable code generation approaches.
V-G Energy Measurement and Hardware Evaluation Gaps
An important research gap identified in this review relates to how energy consumption is measured and evaluated across hardware platforms. As shown in Table XIII, most studies rely on Python-based libraries or operating system–level tools such as perf and Intel RAPL to estimate energy usage. While these tools are convenient and widely used, they provide indirect software-based estimates and may not fully represent actual physical power consumption. Only two studies use external hardware power monitors to measure energy directly, indicating that hardware-based validation is rarely applied.
In addition, most studies evaluate energy consumption on a single hardware platform and do not compare results across different CPUs, GPUs, or embedded devices. Only one study measures energy on two different hardware platforms for comparison. Since energy consumption depends on hardware characteristics such as processor architecture and power management mechanisms, results obtained on a single platform may not generalize to other environments.
Future research should prioritize more rigorous and hardware-aware energy evaluation methodologies in sustainability-focused code generation studies. Specifically, researchers should complement software-based energy estimation tools with external hardware power monitors to validate measurement accuracy and reduce reliance on indirect estimates. In addition, studies should conduct controlled cross-platform experiments across diverse hardware configurations, including CPUs, GPUs, embedded systems, and edge devices, to better understand how energy efficiency varies across computing environments. Such efforts are essential to ensure that conclusions about energy-efficient code generation are reliable, generalizable, and applicable to real-world deployment scenarios.
VI Limitations and Threats to Validity
We finally offer a reflection on the validity threats that are relevant to our study and describe steps we have taken to alleviate those threats. We specifically focus on construct validity, internal validity, and external validity.
VI-A Construct Validity
Construct validity concerns whether the key concepts of this review, sustainable code generation and LLM-based code generation, have been correctly defined and applied during the study selection and analysis process. A challenge in this area is the wide range of terminology used in the literature. For instance, some studies may focus on energy efficiency, carbon emissions, or green software engineering without directly using the term “sustainable code generation.” Likewise, LLM-based systems may be described using terms such as AI-assisted programming, program synthesis, or code completion. Because of these variations, there is a risk of unintentionally including studies that fall outside the intended scope or excluding relevant studies due to differences in wording.
To address this issue, we applied several quality assurance steps as outlined in Section III. While developing the search string, we used a set of manually identified studies as benchmarks to ensure that the search terms properly captured the intended concepts. The search string was refined through multiple iterations to improve coverage and include representative studies. During the selection process, predefined inclusion and exclusion criteria were applied consistently. Any unclear or borderline cases were carefully reviewed and discussed to maintain a shared understanding of the study scope. Similarly, structured data extraction forms were developed and refined through iterative discussion to ensure consistent data collection and classification. These forms were used to systematically record evaluation metrics, model adaptation techniques, and sustainability-related aspects. Although these steps do not eliminate the risk of missing or misinterpreting relevant studies, they help ensure a systematic and consistent identification of the literature included in this review.
VI-B Internal Validity
Internal validity concerns whether the conclusions of this review are properly supported by the evidence presented in the selected studies. In this work, potential threats mainly occur during the study selection and data extraction stages. These risks include selecting studies that do not fully meet the criteria, excluding relevant studies by mistake, and recording incorrect information due to unclear reporting, misunderstanding of the content, or subjective interpretation.
To minimize these risks, we followed clearly defined inclusion and exclusion criteria throughout the screening process. Any uncertain cases were carefully examined and discussed to ensure consistent decisions. We also used structured data extraction forms to systematically collect and classify information such as evaluation metrics, model size categories, prompt engineering techniques, and sustainability aspects. Only information explicitly stated in the papers was recorded. If details were not clearly reported, we did not make assumptions. For instance, when a study used GitHub Copilot without specifying the underlying LLM, it was categorized as “not specified.” Although these steps cannot completely remove all bias, they help maintain a systematic, transparent, and consistent review process.
VI-C External Validity
External validity concerns whether the selected studies represent the existing body of knowledge in sustainable code generation using LLMs. In this review, we included peer-reviewed publications as well as relevant preprints from arXiv, given that this research area is relatively new and rapidly evolving. However, we excluded non-English publications, unpublished student theses, technical blogs, and other forms of grey literature not indexed in the selected databases. As a result, some relevant studies may not have been captured.
In addition, it is possible that certain studies have been accepted for publication but were not yet publicly available at the time of our search. Therefore, our findings reflect the state of the literature based on the selected databases and search timeframe. Despite these limitations, the search string, inclusion and exclusion criteria, and data extraction forms presented in this study provide a transparent framework that can be reused or extended by future researchers to replicate or expand this review.
VII Conclusions
LLMs are increasingly being integrated into software engineering workflows to assist developers with tasks such as code generation, completion, translation, and bug fixing. While significant research has focused on improving the functional correctness and productivity gains of these models, comparatively little attention has been paid to the sustainability of the code they produce. We conduct a systematic literature review to examine whether LLMs can generate sustainable code. We follow a structured process that includes comprehensive database searches and snowballing method, and applying clear inclusion and exclusion criteria in multiple screening stages. For the selected studies, we extract data using predefined forms to ensure consistency and accuracy. Our findings show that research in this area is still limited. Most studies focus on large language models, while small models receive little attention. Energy evaluation is mainly performed at the software level, and only a few studies use external hardware tools for more accurate measurement. In addition, research is largely limited to a small number of programming languages and general-purpose applications. There are no dedicated benchmarks specifically designed to measure the sustainability of LLM-generated code, and sustainability-aware fine-tuning remains an open area for future work. To advance this field, future research should develop dedicated sustainability benchmarks, adopt more accurate energy measurement techniques, and explore sustainability-aware training and fine-tuning strategies. Addressing these gaps is essential to ensure that the growing use of LLMs in software development aligns with long-term environmental sustainability goals.

### v3

@@
-arXiv:2603.00989v2[cs.SE] 03 Mar 2026
+Sustainable Code Generation Using Large Language Models: A Systematic Literature Review
 Abstract
 Large Language Models (LLMs) are increasingly used in software engineering to help developers generate, complete, translate, and fix code.
-These tools improve productivity and speed up the development process.
 However, their growing use has raised concerns about environmental sustainability.
@@
 I Introduction
-Large Language Models (LLMs) are advanced AI systems trained on large corpora of data using transformer-based architectures [131].
+Large Language Models (LLMs) are advanced AI systems trained on large corpora of data using transformer-based architectures [126, 48].
 These models are capable of understanding, reasoning over, and generating human-like language, enabling their use across a wide range of applications[145].
@@
 These capabilities are now embedded in widely used AI-assisted development tools, such as GitHub Copilot [42], CodeGeeX [139], and Amazon CodeWhisperer [7], which leverage state-of-the-art code LLMs to improve developer productivity and streamline the software development process.
-LLMs have achieved tremendous success in Code generation [52, 104]; however, there is a growing concern about the impact of software development on the environment.
+LLMs [90, 47] have achieved tremendous success in Code generation [51, 100]; however, there is a growing concern about the impact of software development on the environment.
 Training and deploying LLMs incurs significant environmental costs, including substantial CO2 emissions and water usage [33].
@@
 Based on this analysis, we identify common patterns, main findings, and existing limitations in the current body of research.
+TABLE I: Research Questions
+RQ Description
+RQ1 How effectively do LLMs produce code that aligns with sustainability principles in practice?
+RQ2 What metrics or parameters are used to evaluate sustainability?
+RQ3 What benchmarks, datasets, and tools are utilized to assess the sustainability of LLM-generated code?
+RQ4 How do prompting strategies or fine-tuning methods influence the sustainability of generated code?
 Overall, we find that research in this domain remains limited.
@@
 The remainder of this paper is organized as follows.
-Section II provides background information on LLMs and code generation.
-Section III describes the research methodology.
-Section IV provides insights into the extracted data, while Section V presents the analysis and results from the extracted data.
-Section VI presents the future research in the area.
-Section VII outlines the validity threats and limitations.
-Finally, Section VIII presents concluding remarks.
+Section II describes the research methodology.
+Section III provides insights into the extracted data, while Section IV presents the analysis and results from the extracted data.
+Section V presents the future research directions.
+Section VI outlines the validity threats and limitations.
+Finally, Section VII presents concluding remarks.
 II Background
@@
 This combination of self-attention and feed-forward layers, together with residual connections and normalization, forms the core building block of the Transformer architecture used in LLMs.
-For more information on transformer architectures and their underlying mechanisms, readers are referred to [131]
+For more information on transformer architectures and their underlying mechanisms, readers are referred to [126].
 Transformer architectures [48, 90] are generally divided into three types: encoder-only, decoder-only, and encoder–decoder [19, 30].
@@
 For a more comprehensive overview of prompt engineering techniques, readers are referred to existing surveys and dedicated studies in this area [77].
-TABLE I: Research Questions
-RQ Description
-RQ1 How effectively do LLMs produce code that aligns with sustainability principles in practice?
-RQ2 What metrics or parameters are used to evaluate sustainability?
-RQ3 What benchmarks, datasets, and tools are utilized to assess the sustainability of LLM-generated code?
-RQ4 How do prompting strategies or fine-tuning methods influence the sustainability of generated code?
 II-B Code Generation
@@
 The following research questions (RQs) were defined to guide the systematic literature review in Table I
-RQ1 is motivated by the need to evaluate how well LLMs can generate code that meets sustainability goals, such as reducing energy consumption, lowering computational overhead, and minimizing resource usage.
-It examines whether sustainability is an inherent property of LLM-generated code or the result of deliberate design choices.
-The scope includes code generated across different application domains and programming languages.
-Addressing RQ1 will help researchers identify which LLMs produce more sustainable code and understand the factors that contribute to this outcome.
-RQ2 aims to identify and categorize the measures used to evaluate the sustainability of LLM-generated code.
-These may include metrics such as energy consumption, carbon footprint, memory usage, execution time, and other efficiency indicators.
-Examining these metrics will show how sustainability is defined and assessed in existing studies.
-The findings will also support the development of standardized evaluation methods, enabling more consistent comparisons across future research.
-RQ3 focuses on identifying the resources used to assess the sustainability of LLM-generated code, including benchmarks, datasets, simulation frameworks, and performance analysis tools.
-It examines whether these resources are specifically designed for sustainability evaluation or adapted from general-purpose code generation benchmarks.
-The findings will help researchers choose appropriate resources for experimentation and highlight gaps where new sustainability-oriented datasets or tools are needed.
-RQ4 examines how model adaptation techniques affect the sustainability of generated code, including prompt engineering, instruction tuning, and domain-specific fine-tuning, etc.
-It focuses on the direct impacts of these techniques on LLM-generated code, such as reduced runtime and lower energy consumption.
-The findings will help establish best practices for designing prompts and fine-tuning methods that enhance sustainability without reducing functional correctness.
+RQ1 evaluates how effectively LLMs generate sustainable code, considering factors such as energy consumption, computational overhead, and resource usage.
+It examines whether sustainability arises inherently from LLMs or from deliberate design choices across different domains and programming languages.
+RQ2 identifies and categorizes the metrics used to assess the sustainability of LLM-generated code, including energy consumption, carbon footprint, memory usage, and execution time.
+This analysis helps clarify how sustainability is defined and supports more standardized evaluation methods.
+RQ3 investigates the benchmarks, datasets, and tools used to evaluate the sustainability of LLM-generated code.
+It assesses whether these resources are specifically designed for sustainability or adapted from general-purpose benchmarks, highlighting gaps for future development.
+RQ4 examines the impact of model adaptation techniques, such as prompt engineering and fine-tuning, on the sustainability of generated code.
+It aims to identify practices that improve efficiency while maintaining correctness.
 III-C Search Strategy
-Guided by the RQs, we adopt an iterative search strategy.
-We repeatedly query digital libraries using Boolean combinations of relevant keywords, and after each iteration, we refine the keywords to improve coverage.
-This process continues until the final search results include all publications from a manually identified benchmark set.
+Guided by the RQs, we adopt an iterative search strategy that queries digital libraries using Boolean combinations of relevant keywords.
+The keywords are refined after each iteration to improve coverage, and the process continues until all manually identified publications are retrieved.
 Figure 1: Primary Study Selection
 III-C1 Search String
-Our search term was structured around two main aspects of the topic.
-The first aspect relates to LLMs, code generation, and associated techniques such as fine-tuning.
-The second aspect focuses on sustainability and energy efficiency.
-Combining these aspects allowed us to capture the broadest possible set of studies relevant to sustainable code generation using LLMs.
-The final search term was formulated by linking keywords from both aspects to ensure that retrieved publications addressed concepts from each domain.
-We refined the search string iteratively by evaluating the results it produced.
-To guide this process, we manually compiled a benchmark set of thirteen (13) relevant research publications from multiple sources, including Google Search, Google Scholar, and arXiv.
-In the first iteration, a comparison between the benchmark set and the search results showed that only 10 out of the 13 publications were retrieved.
-To capture the three missing publications, we added additional terms such as “LLM-generated code,” “energy efficiency,” and “efficient code” to the search string.
-After these additions, the revised search string successfully returned a result set containing the entire benchmark set.
-This final search string was adopted for the remainder of the study and is presented in Table III.
-III-C2 Search Sources
-We searched for relevant publications in five databases: IEEE Xplore, Compendex, Inspec, Scopus, and the ACM Digital Library.
-The retrieved publications were managed using an Excel spreadsheet.
-A detailed breakdown of the search process for each database, along with the corresponding number of publications, is presented in Fig.
+Our search string was structured around two main aspects: (i) LLMs, code generation, and related techniques such as fine-tuning, and (ii) sustainability and energy efficiency.
+Combining these aspects ensured broad coverage of studies on sustainable code generation using LLMs.
+The search string was refined iteratively by evaluating its ability to retrieve relevant studies.
+A benchmark set of thirteen (13) publications was manually compiled from sources such as Google Search, Google Scholar, and arXiv.
+In the initial iteration, 10 out of 13 benchmark studies were retrieved.
+To capture the missing studies, additional terms (e.g., “LLM-generated code,” “energy efficiency,” and “efficient code”) were incorporated.
+The revised search string successfully retrieved the complete benchmark set and was therefore adopted for the remainder of the study.
+The final search string is presented in Table III.
+We searched five databases, IEEE Xplore, Compendex, Inspec, Scopus, and the ACM Digital Library, and managed the retrieved publications using an Excel spreadsheet.
+The search process and publication counts are shown in Fig.
 1.
@@
 (”large language model*” OR ”LLM” OR ”Small Language Model*” OR ”SLM” OR ”GPT” OR ”Codex” OR ”ChatGPT” OR ”code generation” OR ”code generation model*” OR ”AI code assistant” OR ”AI-assisted programming” OR ”automated programming” OR ”code refactoring” OR ”code optimization” OR ”code synthesis” OR ”program synthesis” OR ”code completion” OR ”software language model” OR ”AI programming tool*” OR ”generative AI for code” OR ”code LLM” OR ”AI-based code generation” OR ”prompt engineering” OR ”instruction tuning” OR ”fine-tuning” OR ”zero-shot learning” OR ”few-shot prompting” OR ”chain of thought” OR ”model adaptation” OR ”Llm-generated code”) AND (”sustainable code” OR ”green software” OR ”green software engineering” OR ”energy efficient” OR ”eco-friendly software” OR ”low-energy code” OR ”energy-aware” OR ”carbon footprint” OR ”environmental impact” OR ”code sustainability” OR ”energy-optimized software” OR ”energy-efficient software” OR ”sustainability in software engineering” OR ”sustainable programming” OR ”energy efficiency” OR ”efficient code”)
-TABLE IV: Data Extraction Form
-Category Data Items Description
-General Context of Study Categorized as an academic or industrial setting.
-Code-Related Task The primary coding task addressed (e.g., code generation, completion, etc.
-Coding Application The application domain of generated code (e.g., blockchain, scientific computing, etc).
-Model Evaluated Models and Size Categories Lists and classify the LLM(s) used in the study, along with their parameter sizes.
-Model Openness Specifies whether the models are open or closed-source, or a combination of both.
-Model Domain Specialization Indicates whether the model is general-purpose, code-specific, or a combination of both.
-Evaluation Evaluation Metrics Metrics used to assess performance and sustainability.
-Custom Evaluation Metric Indicates whether the study introduces a novel evaluation metric.
-Libraries / Tools for Measurement Tools or libraries used for metric collection.
-Hardware Configuration Description of the hardware and experimental setup used for code execution.
-Experimental Configuration Details regarding experiment repetitions and inference parameters (e.g., temperature, top-p, maximum tokens, batch size, precision).
-Open-Source Availability Availability of code, prompts, datasets, or other resources to support reproducibility.
-Benchmark Usage Indicates whether established benchmarks are used for model evaluation.
-Model Adaptation Fine-Tuning Details If the study is performing sustainability-aware fine-tuning.
-Use of Prompt Engineering Specifies whether prompt engineering strategies are applied to generate sustainable code.
-III-D Inclusion and Exclusion Criteria
-To ensure that only the most relevant and high-quality studies were considered in this systematic literature review, we applied a predefined set of inclusion and exclusion criteria, as summarized in Table II.
-In order to qualify as a primary study, the research work must meet all inclusion criteria and must not meet any exclusion criteria.
-We included studies that used or evaluated LLMs for code generation, addressed sustainability aspects such as energy efficiency or resource usage, and provided empirical evaluation results.
-Studies were excluded if they did not involve an LLM, lacked sustainability analysis, focused on unrelated tasks, or had no accessible full text in English.
-The inclusion and exclusion criteria were applied at multiple stages of the literature review process, including the initial filtering of primary studies and the full-text screening phase.
-They were also employed during the snowballing process, which resulted in the addition of 4 more primary studies.
-Although several studies claim efficiency improvements, many of them do not provide empirical evaluation of energy consumption or energy efficiency.
-Since our objective is to examine the sustainability of LLM-generated code based on measurable evidence, we refined our exclusion criteria to remove studies that do not include any form of energy consumption analysis.
-As a result, only studies that provide explicit energy-related evaluation were considered in the final selection.
-This refinement reduced the number of eligible studies to 19.
-Moreover, to ensure comprehensive coverage of this emerging research area, we also included relevant pre-prints from arXiv, as sustainability-focused research on LLM-generated code is still developing and may not yet be fully represented in peer-reviewed venues.
 Study Models Evaluated Model Size Category
@@
 [26] GitHub Copilot (StarCoder family back-end models: StarCoder 15.5 B, StarCoder2-7 B, StarCoder2-15 B) Mixed (Small + Medium)
-[60] Llama 3.2 (3 B) and OPT (2.7 B) Small
 [24] GPT-4 Large
@@
 [9] ChatGPT (175B), GPT-4 (1.76T), DeepSeek Coder Instruct (33B), Speechless CodeLlama v2.0 (34B), CodeMillennials (34B), WizardCoder v1.1 (33B) Large
-TABLE V: Overview of the evaluated models and their parameter-scale categories.
+TABLE IV: Overview of the evaluated models and their parameter-scale categories.
+III-D Inclusion and Exclusion Criteria
+To ensure that only relevant and high-quality studies were included in this systematic literature review, we applied a predefined set of inclusion and exclusion criteria, summarized in Table II.
+We included studies that used or evaluated LLMs for code generation, addressed sustainability aspects such as energy efficiency or resource usage, and reported empirical results.
+Studies were excluded if they did not involve LLMs, lacked sustainability analysis, focused on unrelated tasks, or did not have accessible full text in English.
+The criteria were applied at multiple stages, including initial filtering, full-text screening, and snowballing, which resulted in the addition of 3 more primary studies.
+Although several studies claim efficiency improvements, many do not provide empirical evaluation of energy consumption.
+Since our objective is to assess the sustainability of LLM-generated code based on measurable evidence, we refined the exclusion criteria to remove studies lacking energy-related analysis.
+As a result, only studies with explicit energy evaluation were included, reducing the final set to 18 studies.
+To ensure comprehensive coverage of this emerging area, we also included relevant preprints from arXiv, as sustainability-focused research on LLM-generated code is still evolving and may not yet be fully represented in peer-reviewed venues.
 III-E Study Selection Procedure
@@
 In the second iteration, Fleiss’ Kappa inter-rater agreement increased to 0.833, indicating that the researchers had reached a sufficient and consistent understanding of the inclusion criteria.
-We then applied the refined inclusion criteria independently to the entire set of retrieved publications, identifying 170 studies for full-text review.
-To maintain consistency, we used the same pilot process with Fleiss’ Kappa measurement during this stage.
-After carefully reading the full texts, we found that only 15 studies fully met the refined criteria and were included as primary studies.
-We chose not to apply additional quality assessment filters, as the final pool of studies was already small and the research area is still emerging.
-This approach allowed us to remain as inclusive as possible while ensuring relevance to our research focus.
-We conducted a snowballing [66] search to identify additional relevant primary studies.
-Snowballing involves reviewing the reference lists of papers (backward snowballing) and the papers that cite them (forward snowballing) to find further studies.
-In this study, both forward and backward snowballing were applied to the 15 retrieved primary studies.
-Google Scholar was used to obtain the citations for each paper, and the same selection criteria were applied to both their references and citations.
-This process yielded 4 additional relevant papers, bringing the final pool of primary studies to 19.
-Figure 1 shows how the initial 2,015 publications retrieved from the databases were reduced to these 19 primary studies, which were then used for data extraction.
-III-F Data Extraction Form
+We then applied the refined inclusion criteria to all retrieved publications, identifying 170 studies for full-text review.
+To ensure consistency, the same pilot process with Fleiss’ Kappa measurement was used at this stage.
+After full-text screening, only 15 studies met the criteria and were included as primary studies.
+No additional quality assessment filters were applied, as the final pool was already small and the research area is still emerging.
+This approach ensured inclusiveness while maintaining relevance to our research focus.
+We conducted a snowballing [64] search to identify additional primary studies.
+This involved reviewing both reference lists (backward snowballing) and citing papers (forward snowballing).
+Both approaches were applied to the 15 selected studies, using Google Scholar to retrieve citations.
+The same selection criteria were applied to all references and citations.
+This process identified 3 additional studies, bringing the final set to 18 primary studies.
+Fig.
+1 illustrates how the initial 2,015 publications were reduced to these 18 studies used for data extraction.
+TABLE V: Data Extraction Form
+Category Data Items Description
+General Context of Study Categorized as an academic or industrial setting.
+Code-Related Task The primary coding task addressed (e.g., code generation, completion, etc.)
+Coding Application The application domain of generated code (e.g., blockchain, scientific computing, etc).
+Model Evaluated Models and Size Categories Lists and classify the LLM(s) used in the study, along with their parameter sizes.
+Model Openness Specifies whether the models are open or closed-source, or a combination of both.
+Model Domain Specialization Indicates whether the model is general-purpose, code-specific, or a combination of both.
+Evaluation Evaluation Metrics Metrics used to assess performance and sustainability.
+Custom Evaluation Metric Indicates whether the study introduces a novel evaluation metric.
+Libraries / Tools for Measurement Tools or libraries used for metric collection.
+Hardware Configuration Description of the hardware and experimental setup used for code execution.
+Experimental Configuration Details regarding experiment repetitions and inference parameters (e.g., temperature, top-p, maximum tokens, batch size, precision).
+Open-Source Availability Availability of code, prompts, datasets, or other resources to support reproducibility.
+Benchmark Usage Indicates whether established benchmarks are used for model evaluation.
+Model Adaptation Fine-Tuning Details Indicates whether the study is performing sustainability-aware fine-tuning.
+Use of Prompt Engineering Specifies whether prompt engineering strategies are applied to generate sustainable code.
+IV Data Extraction Form
 We created a structured data extraction form to collect information from each primary study consistently and accurately, focusing on our research questions (RQ1–RQ4).
 First, three researchers prepared a draft extraction form based on a small sample of studies.
-We then met to combine these drafts into the final extraction form shown in Table IV.
+We then met to combine these drafts into the final extraction form shown in Table V.
 The categories in the form were designed to capture the main differences found in research on sustainable code generation using LLMs.
@@
 The extracted data offers a clear framework for analysis, enabling the comparison and summarization of findings across all studies. The results are presented in the next section.
-IV Data Extraction
-We present the analysis of data collected from the final pool of 19 primary studies.
-This section outlines the extracted data, as summarized in Table IV, providing a structured basis for analysis and enabling comparisons across all studies.
-The information is then used in Section LABEL:subsec:results_rqs to synthesize findings and address RQ1–RQ4.
-IV-A General
-IV-A1 Context of Study
-We first categorized the selected studies based on the environment in which the research was undertaken, distinguishing between academic and industrial settings.
-A study was labelled as academic if all contributing authors were affiliated with universities or research institutions.
-A study was labelled as industrial if the research involved participation from industry, identified by the presence of at least one author with an industrial affiliation.
-The distribution of studies across these contexts is summarized in Table VI.
-Out of the 19 studies, 14 were conducted in an academic setting, while only 5 were performed in an industrial setting.
+V Data Extraction
+We present the analysis of data from the final set of 18 primary studies.
+This section outlines the extracted data, as summarized in Table V, providing a structured basis for analysis and enabling comparisons across all studies.
+This information is used in Section VI to synthesize findings and address RQ1–RQ4.
+V-A General
+V-A1 Context of Study
+We first categorized the selected studies based on the research environment, distinguishing between academic and industrial settings.
+A study was labelled as academic if all authors were affiliated with universities or research institutions, and as industrial if at least one author had an industry affiliation.
+The distribution is summarized in Table.
+VI.
+Of the 18 studies, 13 were academic and 5 were industrial.
 Infrastructure Type Primary Studies Total
-Academic [96], [110], [127], [33], [22], [27], [60], [25], [105], [101], [97], [116], [29], [10] 14
+Academic [92], [105], [122], [32], [21], [26], [24], [101], [97], [93], [111], [28], [9] 13
 Industrial [125], [124], [62], [10], [11] 5
 TABLE VI: Context of study
-IV-A2 Code-Related Task
+V-A2 Code-Related Task
 We categorized the selected studies based on the code-related tasks they addressed.
@@
 As several studies addressed more than one task (e.g., combining code generation with optimization or refactoring), a multi-label classification approach was adopted; therefore, a single study may appear in multiple task categories.
-This approach enables a more accurate representation of the scope of each study and avoids oversimplifying contributions that span multiple tasks.
 Infrastructure Type Primary Studies
-Code Generation [33], [127], [22],[27], [60], [25], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10]
+Code Generation [32], [122], [21],[26], [24], [125], [124], [62], [97], [10], [93], [111], [28], [11], [9]
 Code Optimization [92], [101], [97], [111]
@@
 In contrast, relatively fewer studies address code optimization, code completion, or code refactoring, suggesting that these tasks have received limited attention from an energy measurement perspective.
-IV-A3 Code Application Domain
+V-A3 Code Application Domain
 To better understand the application contexts of the selected studies, we categorized them based on the type of software and workloads they target.
@@
 Infrastructure Type Primary Studies Total
-General-Purpose Software Development [96], [110], [127], [22], [27], [60], [25], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10] 17
+General-Purpose Software Development [92], [105], [122], [21], [26], [24], [125], [124], [62], [97], [10], [93], [111], [28], [11], [9] 16
 Scientific and Engineering Applications [101], [32] 2
@@
 In [101], the authors evaluate the effectiveness of LLMs in reducing the environmental footprint of real-world MATLAB projects.
-IV-B Model
-IV-B1 Evaluated Models and Model Size Categories
-We first present the list of models considered in each study in Table V.
-The models are listed as reported by the original authors to accurately reflect the experimental settings used in each work.
-One study does not explicitly report the underlying models employed and is therefore marked as Not Specified.
-To enable consistent comparison across studies, we also categorize the evaluated language models based on their parameter count into three size categories: small, medium, and large.
-Small models have at most 7 billion parameters ($\leq 7B$), medium models have between 8 and 30 billion parameters, and large models exceed 30 billion parameters ($\geq 30B$).
+V-B Model
+V-B1 Evaluated Models and Model Size Categories
+We first present the list of models considered in each study in Table IV. The models are listed as reported by the original authors to accurately reflect the experimental settings used in each work.
+To enable consistent comparison, we categorize models by parameter count into three groups: small ($\leq 7B$), medium (8–30B), and large ($\geq 30B$).
 Some studies evaluate models spanning multiple size categories and are therefore labelled as mixed-scale to accurately reflect their evaluation scope.
@@
 This classification enables a clear and reproducible analysis of model usage trends while accounting for differences in computational scale.
-IV-B2 Model Openness
+V-B2 Model Openness
 To further characterize the experimental settings of the selected studies, we categorized the evaluated models based on their availability and licensing.
@@
 Category Primary Studies Total
-Open-Source [110], [22], [27], [60], [29], [12] 6
+Open-Source [105], [21], [26], [28], [11] 5
 Closed-Source [92], [24], [125], [124], [97], [111] 6
@@
 Table IX summarizes the categorization of the selected studies based on model openness.
-It can be seen that the studies are evenly distributed across open-source, closed-source, and mixed model categories, with six studies in each group.
+It can be seen that the studies are distributed across open-source, closed-source, and mixed model categories, with five studies using open-source models and six studies each in the closed-source and mixed categories.
 In addition, although one study [32] discusses the use of proprietary models, it does not clearly specify which models are evaluated and is therefore categorized as not specified.
-IV-B3 Model Domain Specialization
+V-B3 Model Domain Specialization
 To characterize the domain focus of the evaluated models, we classify them as general-purpose, code-specialized, or mixed (general and code-specialized) language models.
@@
 Studies that evaluate both general-purpose and code-specialized models are classified as mixed.
-Studies that do not report the domain specialization of the evaluated models are labelled as not specified.
+Studies that do not report the domain specialization of the evaluated models are labeled as not specified.
 Category Primary Studies Total
-General-Purpose LLMs [96], [110], [60], [25], [105] 5
+General-Purpose LLMs [92], [105], [24], [101] 4
 Code-Specialized LLMs [21], [26], [125], [97], [111], [28], [11] 7
@@
 The distribution of studies across these domains is presented in Table X.
-It can be seen that only five studies focus exclusively on general-purpose language models, while the remaining studies evaluate either code-specialized models or include at least one code-specialized model in a mixed evaluation setting.
-IV-C Evaluation
-IV-C1 Evaluation Metrics
+It can be seen that only four studies focus exclusively on general-purpose language models, while the remaining studies evaluate either code-specialized models or include at least one code-specialized model in a mixed evaluation setting.
+V-C Evaluation
+V-C1 Evaluation Metrics
 We categorize the evaluation metrics considered in the primary studies into six groups: energy, performance, correctness, memory, code quality, and miscellaneous metrics.
@@
 For each metric, the table lists the primary studies in which it is used, providing a structured overview of evaluation practices in the literature.
-Table XI summarizes the evaluation metrics reported across the selected studies.
-It can be observed that most studies focus on energy consumption and performance metrics, particularly latency and runtime.
-Fewer studies evaluate functional correctness and memory usage, and only a small number consider code quality.
+Study Custom Evaluation Metrics
+[122] RuntimeRatio, EnergyRatio
+[32] Energy-reduction@k
+[125] Embodied Energy, Operational Energy, Carbon Intensity, Embodied Carbon Emissions, Operational Carbon Emissions
+[124] Green Capacity (GC)
+TABLE XI: Studies proposing custom sustainability-oriented evaluation metrics.
 Category Evaluation Metric Primary Studies
-Energy Energy consumption [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10]
+Energy Energy consumption [92], [105], [21], [26], [24], [101], [125], [124], [62], [97], [10], [93], [111], [28], [11], [9]
 Power consumption [26], [11]
@@
 Carbon footprint [24]
-Performance Latency/Runtime [96], [110], [22], [27], [60], [25], [105], [130], [129], [64], [101], [11], [97], [116], [12]
+Performance Latency/Runtime [92], [105], [21], [26], [24], [101], [125], [124], [62], [97], [10], [93], [111], [11]
 Runtime Ratio (%) [122]
-Throughput (tokens/s) [60], [97]
+Throughput (tokens/s) [93]
 CPU cycles [93]
@@
 Code Quality Exact Match [105]
-CodeBLEU, ROUGE-L [60]
-Syntax & DataFlow score [60]
 AST similarity (%) [21]
@@
 Completed generations [26]
-TABLE XI: Categorization of evaluation metrics reported in the selected studies.
-IV-C2 Custom Evaluation Metrics
+TABLE XII: Categorization of evaluation metrics reported in the selected studies.
+Table XII summarizes the evaluation metrics reported across the selected studies.
+It can be observed that most studies focus on energy consumption and performance metrics, particularly latency and runtime.
+Fewer studies evaluate functional correctness and memory usage, and only a small number consider code quality.
+V-C2 Custom Evaluation Metrics
 In addition to standard evaluation metrics, a small number of studies introduce custom sustainability-oriented metrics.
 These metrics are designed by the authors to capture specific aspects of energy efficiency, carbon impact, or sustainability that are not covered by conventional measures.
-Table XII presents the studies that propose such custom evaluation metrics and lists the corresponding metrics introduced in each work.
+Table XI presents the studies that propose such custom evaluation metrics and lists the corresponding metrics introduced in each work.
 Studies that do not propose any custom evaluation metrics are omitted from this table.
-Study Custom Evaluation Metrics
-[127] RuntimeRatio, EnergyRatio
-[33] Energy-reduction@k
-[130] Embodied Energy, Operational Energy, Carbon Intensity, Embodied Carbon Emissions, Operational Carbon Emissions
-[129] Green Capacity (GC)
-TABLE XII: Studies proposing custom sustainability-oriented evaluation metrics.
-RuntimeRatio and EnergyRatio compare the runtime and energy consumption of model-generated code against canonical human-written solutions, providing a relative efficiency measure.
-Energy-reduction@k estimates the expected energy savings when generating multiple candidate solutions and selecting the most energy-efficient valid one, while accounting for both performance improvement and correctness probability.
-Similarly, Green Capacity (GC) measures the overall sustainability improvement of optimized code across multiple performance dimensions, filtering out invalid or degraded solutions.
-Embodied Energy and Embodied Carbon capture the environmental impact associated with code generation or development infrastructure, whereas Operational Energy and Operational Carbon measure emissions during actual code execution.
-Carbon Intensity (CI) quantifies the amount of carbon emissions per unit of electricity consumed.
-IV-C3 Libraries / Tools for Measurement
+V-C3 Libraries / Tools for Measurement
 To understand how energy is measured in the selected studies, Table XIII presents the tools and libraries used for measuring energy consumption and related system metrics.
@@
 Perf[96] [21], [26], [124], [62], [111]
-PyNVML[91] [33], [60]
+PyNVML[87] [32]
 Nvidia-smi[88] [26]
@@
 TABLE XIII: Tools and libraries used for energy measurement
-IV-C4 Hardware Configuration
+V-C4 Hardware Configuration
 Table XIV summarizes the hardware environments used for evaluation in the selected studies.
@@
 Finally, not specified indicates studies that do not explicitly report the hardware used for evaluation, which limits reproducibility and comparability.
-Evaluation Hardware Category Representative Configurations
-On-premise (local machine/server) [110], [127], [22], [27], [60], [129], [64], [116], [10]
+Evaluation Hardware Category Primary Studies
+On-premise (local machine/server) [105], [122], [21], [26], [124], [62], [111], [9]
 Cloud [97], [10], [93], [11]
@@
 TABLE XIV: Categorization of hardware configurations used for evaluation in the selected studies.
-IV-C5 Experimental Configuration
+V-C5 Experimental Configuration
 Table XV summarizes the experimental configuration, including the number of runs, decoding strategy [134], temperature [120], maximum token limits, and top-p [85] and top-k [138] values.
@@
 [26] 5 - - - -
-[60] - - 0.6-0.92 15 -
 [24] 3 - - - -
@@
 TABLE XV: Decoding and inference settings reported in the selected studies.
-IV-C6 Open-Source Availability
+V-C6 Open-Source Availability
 To assess reproducibility and transparency, we examine the open-source availability of the selected studies.
@@
 Table XVI summarizes the open-source availability of the reviewed studies, indicating whether implementation artifacts are publicly released.
-Category Count
-Open-source available [96], [110], [33], [22], [27], [60], [25], [105], [64], [97], [116], [29], [10]
+Category Primary Studies
+Open-source available [92], [105], [32], [21], [26], [24], [101], [62], [93], [111], [28], [9]
 Not open-source [122], [125], [124], [97], [10], [11]
 TABLE XVI: Summary of open-source availability
-IV-C7 Benchmark Comparison and Usage
-We also categorize the evaluation data based on the type of benchmark or dataset used and the programming language considered.
-Table XVII summarizes the benchmarks and datasets used in the selected studies, along with programming language information when explicitly reported.
+V-C7 Benchmark Comparison and Usage
+We categorize the evaluation data based on the type of benchmark or dataset used and the programming language considered.
+Table XVII summarizes the benchmarks and datasets, along with programming language information when explicitly reported.
 Programming languages are reported only when explicitly specified by the authors or when the language choice is central to the evaluation; for standardized benchmarks, the programming language is implicit and therefore omitted.
-It can be observed that most studies primarily focus on Python, reflecting its widespread use in code generation research.
-In addition, not all studies rely on standardized benchmarks, as several works evaluate models using diverse or custom data sources, indicating variability in evaluation practices across the literature.
+Most studies primarily focus on Python, reflecting its widespread use in code generation research.
+Additionally, not all studies use standardized benchmarks, with several relying on custom data sources, indicating variability in evaluation practices.
 Study Benchmark / Dataset Source Programming Language
@@
 [26] CLI Connect 4 Game[119] Java
-[60] JavaCorpus[6], PY150[106] Java, Python
 [24] Codeforces[25] Python
@@
 TABLE XVII: Benchmarks and programming languages used in the selected studies.
-IV-D Model Adaptation
-IV-D1 Fine-Tuning
-We also categorize the studies based on the use of fine-tuning, that is, whether a study adapts the model parameters to generate more sustainable code.
-Surprisingly, only one study [60] employs fine-tuning for this purpose.
-This study proposes a fine-tuning approach for code generation models using a weighted aggregated loss [77], which enables models to perform dynamic early exits at different layers without requiring additional language model heads.
-In addition, it introduces GREEN-CODE, a reinforcement learning–based [71] framework that improves energy efficiency in code generation by dynamically selecting early exit points during inference.
-IV-D2 Prompt Engineering
-Table XVIII summarizes whether the selected studies use prompt engineering techniques.
-A study is marked as “Yes” if it explicitly applies methods such as zero-shot, few-shot, chain-of-thought, iterative prompting, or prompt variations to influence code generation.
-Studies that use default model settings without modifying prompts are marked as “No.” It can be observed that several studies rely on prompt-based strategies, while others evaluate models without applying specific prompt engineering techniques.
+V-D Model Adaptation
+V-D1 Fine-Tuning
+We categorize the studies based on the use of fine-tuning, i.e., whether model parameters are adapted to generate more sustainable code.
+Notably, none of the reviewed studies employ fine-tuning specifically for improving sustainability.
+This highlights an open research direction in leveraging fine-tuning to guide LLMs toward more energy-efficient and resource-aware code generation.
+V-D2 Prompt Engineering
+Table XVIII summarizes the use of prompt engineering techniques in the selected studies.
+A study is marked as “Yes” if it explicitly applies methods such as zero-shot, few-shot, chain-of-thought, or prompt variations, and as “No” if default prompts are used.
+Several studies employ prompt-based strategies, while others evaluate models without specific prompt engineering.
+VI Results and Synthesis
+VI-A RQ1: How effectively do LLMs produce code that aligns with sustainability principles in practice?
+The purpose of this RQ is to evaluate whether LLMs can generate sustainable code in practice by examining the consistency of efficiency improvements across tasks and models, as well as key trade-offs such as energy efficiency and functional correctness.
+Although sustainability is often an explicit objective in the reviewed studies, the empirical evidence does not consistently show that LLM-generated code outperforms human baselines.
+Among the selected studies, 7 studies reported that LLM-generated code does not outperform human baselines in sustainability-oriented comparisons ( [124],[122],[24],[101],[62],[28],[9]).
+Within this group, 2 studies explicitly concluded that human-written code performs better than LLM-generated code ( [124, 28]).
+In addition, 3 studies characterized the outcome as inconsistent, indicating that any observed improvement depends on the task, or benchmark ([21, 125, 10]).
+The remaining 8 studies did not evaluate a direct human-versus-LLM comparison for sustainability outcomes ([92],[105], [32],[26],[97],[93],[111],[11].
+Beyond the human-versus-LLM comparison, the reported sustainability outcomes are frequently context-dependent.
+Specifically, 14 studies indicate that efficiency gains are task-dependent, varying across benchmarks, problem types, or code characteristics ([124], [92], [122], [32], [21], [24], [101], [62], [97], [10], [93], [111] , [28], [11] ).
+Moreover, 11 studies report model-dependent outcomes, suggesting that sustainability performance can differ across LLMs and code assistants even under similar evaluation settings ([124], [21], [26], [101], [62], [10], [93], [11] , [9]) and one additional study reporting model sensitivity.
+Trade-offs are also a dominant theme in this literature with 16 studies explicitly discuss trade-offs ([124], [92], [105], [122], [32], [26], [24], [101], [125], [62], [97], [10], [93], [111] , [28], [9]).
+Within this group, 14 studies explicitly analyze the energy and accuracy trade-off ([124], [92], [105], [122], [32], [24], [101], [125], [62], [97], [10], [93], [111] , [9]), indicating that many studies treat sustainability as a multi-objective problem rather than a single-metric improvement.
+The remaining two studies discuss trade-offs but do not directly frame them as energy-versus-correctness ([26], [28]).
+Only two studies do not report an explicit trade-off discussion ([21], [11]).
+Overall, this pattern suggests that sustainability gains are frequently evaluated alongside potential costs such as correctness, and the results depend on how studies balance energy savings and correctness.
 Prompt Engineering Primary Studies
 Yes [92], [105], [32], [21], [26], [97], [93], [28], [11], [9]
-No [127], [60], [25], [105], [130], [129], [64], [11], [116]
-TABLE XVIII: Use of prompt engineering techniques in the selected studies.
-V Future Research Opportunities
+No [122], [24], [101], [125], [124], [62], [10], [111]
+TABLE XVIII: Use of prompt engineering techniques
+Most studies evaluate sustainability using a single final output per prompt/run, rather than generating multiple candidate solutions.
+Specifically, 12 studies report single-solution evaluation ([124], [92], [105], [122], [21], [24], [101], [125], [10], [111], [28], [11]).
+Only two studies employ Pass@k multi-sampling ([32], [62]), and among these, only one study explicitly reports an energy@k perspective ([32]).
+In addition, two studies generate multiple candidates through re-prompting without reporting Pass@k or energy@k metrics ([97], [9]), while one study does not clearly specify whether multi-sampling is used ([26]).
+The energy overhead of generating multiple candidates is also largely under-reported.
+Only two studies explicitly state that producing additional candidates increases energy consumption ([32], [24]), and one study notes this effect qualitatively by indicating that additional refinement steps may increase overall cost, such as energy, latency ([92]).
+The remaining 15 studies do not quantify or report the energy impact of multi-candidate generation ([124], [105], [122], [21], [26], [101], [125], [62], [97], [10], [93], [111], [28], [11], [9]).
+Overall, these distributions indicate that, although multi-sampling and iterative prompting are common in practice, most studies assess sustainability using single-output settings and do not measure the additional energy overhead introduced by candidate sampling, limiting conclusions about end-to-end sustainability under realistic usage patterns.
+Furthermore, the relationship between sustainability and correctness is mixed across the selected studies.
+Specifically, 8 studies report that energy-efficient outputs are also correct ([92], [105], [32], [62], [97], [93], [111], [9]), whereas 6 studies report that energy-efficient outputs are not consistently correct ([122], [24], [101], [125], [28], [11]); 2 studies report mixed outcomes ([124], [10]), and 2 studies do not evaluate this aspect ([21], [26]).
+Evidence regarding model-size effects remains limited.
+Only 2 studies report that larger models generate more energy-efficient code ([122], [10]), whereas 4 studies report contrary findings ([92], [101], [62], [11] ) and the remaining 13 studies do not evaluate model-size effects.
+Notably, none of the studies discuss energy-per-token, indicating a consistent reporting gap.
+Regarding small models, 2 studies report that small models achieve roughly the same efficiency as large models ([26], [10]).
+In contrast, 5 studies indicate that small models do not achieve comparable efficiency ([124], [92], [122], [101], [62]).
+Moreover, 2 studies report inconsistent findings, where the outcome depends on the task or evaluation setting ([21], [11] ).
+The remaining 9 studies do not evaluate this aspect ([105], [32], [24], [125], [97], [93], [111] , [28], [9]).
+Finally, only 7 studies explicitly compare small and large models ([122], [21], [26], [62], [10], [93], [9]), whereas the remaining 11 studies do not evaluate this comparison ([124], [92], [105], [32], [24], [101], [125], [97], [111] , [28], [11] ).
+This limited coverage indicates that conclusions about sustainability effects across model scales remain constrained by the relatively small number of comparative studies.
+VI-B RQ2: What metrics or parameters are used to evaluate sustainability?
+This research question investigates how sustainability is evaluated in the literature by identifying the most commonly used metrics such as energy consumption, execution time, memory usage, carbon footprint, etc.
+Across the selected studies, sustainability evaluation is dominated by a small set of quantitative metrics.
+In all the studies, the sustainability of the generated code is measured using the metric energy consumption in 16 studies.
+Other aspects of sustainability are evaluated using the metric power consumption in two studies and the carbon footprint in one study.
+Two studies used custom evaluation metrics to measure the energy of the generated code, i.e., EnergyRatio, and Energy-Reduction@k.
+These studies are shown in Table.
+XII.
+In addition to energy consumption, 9 studies also reported memory-related metrics to account for the sustainability of the generated code.
+This suggests that sustainability has been measured using different evaluation metrics, however some metrics are more commonly used, such as energy consumption, compared to carbon footprint.
+Finally, 4 studies introduce custom metrics ([124], [122], [32], [125]), typically motivated by capturing different dimensions of sustainability.
+RuntimeRatio and EnergyRatio compare the runtime and energy consumption of model-generated code against canonical human-written solutions, providing a relative efficiency measure.
+Energy-reduction@k estimates the expected energy savings when generating multiple candidate solutions and selecting the most energy-efficient valid one, while accounting for both performance improvement and correctness probability.
+Similarly, Green Capacity measures the overall sustainability improvement of optimized code across multiple performance dimensions, filtering out invalid or degraded solutions.
+Embodied Energy and Embodied Carbon capture the environmental impact associated with code generation or development infrastructure, whereas Operational Energy and Operational Carbon measure emissions during actual code execution.
+Carbon Intensity (CI) quantifies the amount of carbon emissions per unit of electricity consumed.
+VI-C RQ3: What benchmarks, datasets, and tools are utilized to assess the sustainability of LLM-generated code?
+The analysis of the selected studies shows that researchers rely on a diverse set of benchmarks and datasets to evaluate LLM-generated code, although several common sources emerge.
+A significant portion of the studies use programming challenge platforms, such as LeetCode, Codeforces, and HackerRank, where authors typically collect or design custom subsets of problems or code solutions from these websites rather than relying on a fixed benchmark dataset [122, 21, 24, 124, 10, 28, 11].
+These platforms provide algorithmic programming tasks that are widely used to assess the correctness and efficiency of generated code.
+In addition to these sources, several studies employ established code generation benchmarks, including CodeXGLUE, HumanEval, EvoEval, and EffiBench, which are specifically designed for evaluating LLM-based code generation and software engineering tasks [105, 62, 111, 9].
+Other works use large-scale code corpora and repositories, such as GitHub repositories, to analyze real-world code [101].
+Some studies further rely on domain-specific benchmark suites, including HeCBench, XSBench, miniMDock, SciMark2, and DaCapoBench, which are commonly used in scientific computing and performance evaluation contexts [32, 93].
+A few studies also introduce custom datasets or applications, such as the Energy-Language dataset or a CLI-based Connect-4 game [92, 26], while some studies do not clearly specify the dataset used [125, 97].
+With respect to programming languages, the experiments predominantly focus on Python, followed by C++ and Java, with occasional inclusion of JavaScript and MATLAB.
+Several programming languages, such as Go and Solidity, are not represented in the reviewed studies and therefore remain unexplored in the context of sustainable code generation.
+Overall, the results indicate that most studies rely on custom datasets derived from programming challenge platforms or other problem sources, rather than consistently using standardized benchmarks for evaluating the sustainability of generated code.
+The reviewed studies employ a variety of tools and libraries to measure the energy consumption of LLM-generated code, with most relying on software-based energy estimation frameworks.
+Commonly used tools include Intel RAPL, CodeCarbon, PyJoules, Perf, and PyNVML, which estimate energy consumption by accessing CPU or GPU performance counters and system-level statistics [92, 105, 122, 21, 24, 124, 111].
+Additional system monitoring utilities such as Nvidia-smi, Rocm-smi, and Turbostat are also used to obtain hardware utilization and power-related metrics [26, 93].
+However, very limited number of studies perform direct hardware-level energy measurements using external power monitoring devices.
+In particular, [28, 9] employ a Monsoon Power Monitor, specifically on Raspberry Pi platforms.
+VI-D RQ4: How do prompting strategies or fine-tuning methods influence the sustainability of generated code?
+Among the 18 primary studies, prompt engineering techniques explicitly targeting energy efficiency or sustainability are investigated in 10 studies ([92], [105], [32], [21], [26], [97], [93], [28], [11], [9]).
+The remaining 8 studies do not vary prompts for sustainability objectives ([122], [24], [101], [125], [124], [62], [10], [111]).
+Prompting strategies range from basic methods like zero-shot (only task description) and few-shot (with examples), to more structured approaches such as role-based prompting (assigning the model a specific role) and energy-specialised prompting (explicitly asking for efficient or low-resource outputs).
+More advanced techniques, including Chain-of-Thought (CoT) reasoning and iterative feedback loops, are also used to improve both performance and energy efficiency.
+Together, these prompting methods form a progression from simple instructions to more controlled and optimisation-focused designs.
+Zero-shot prompting does not consistently reduce the energy consumption of generated code.
+Some studies report results similar to or even worse than baseline human-written or unoptimized code [124, 122].
+In contrast, Chain-of-Thought (CoT) prompting shows more consistent improvements, particularly for small and medium-sized models.
+For example, [11] demonstrates that CoT outperforms baseline approaches for models such as Qwen2.5-Coder-3B and StableCode-3B, although it provides limited benefits for larger models like CodeLlama-7B and Phi-3-Mini-4K.
+Similarly, [9] reports energy savings with CoT, though the results vary across hardware platforms.
+Explicit energy-aware prompting produces mixed results across studies.
+Some works show measurable reductions in energy consumption [101, 97, 9], with one study reporting up to a 36% reduction when combined with iterative feedback [97].
+However, other studies find that such prompts can lead to neutral or inconsistent outcomes compared to baseline prompting [122, 21].
+Iterative prompting strategies, which incorporate self-correcting feedback loops and LLM-as-a-judge mechanisms, show the strongest and most consistent improvements in the studies that adopt them.
+These approaches achieve significant reductions in energy consumption while maintaining functional correctness, although they require multiple generation steps [32, 97].
+Overall, no single prompting strategy consistently outperforms others across all settings.
+The effectiveness of a given approach depends on factors such as the model, the nature of the task, and the execution environment.
+In general, more structured prompting techniques, such as Chain-of-Thought, tend to perform better than simpler approaches like zero-shot in certain scenarios, while specialized and iterative prompting methods often provide additional benefits in practical, real-world applications.
+VII Future Research Opportunities
 We now present general observations and insights from our systematic review on sustainable code generation using LLMs, highlighting key research gaps and directions for future work.
-V-A Expanding Sustainability Beyond Code Generation
+VII-A Expanding Sustainability Beyond Code Generation
 One notable observation is the need to extend sustainability evaluation beyond code generation tasks.
@@
 Since software sustainability depends not only on generating code but also on improving and maintaining it over time, these areas represent significant and currently underexplored opportunities for advancing energy-aware LLM research.
-V-B Underexplored Application Domains
+VII-B Underexplored Application Domains
 An important gap identified in this review is the limited exploration of sustainability-aware code generation across diverse application domains.
@@
 For example, embedded and IoT systems operate under strict energy constraints, low-level software can significantly influence hardware efficiency, and cloud-based services contribute substantially to large-scale energy consumption.
-The limited attention to these areas highlights the need for future studies to evaluate LLM-generated code in more realistic, energy-sensitive, and deployment-oriented environments.
-V-C Expanding Sustainability Evaluation Across Programming Languages
+The limited attention to these areas highlights the need for future studies to evaluate LLM-generated code in more realistic, and deployment-oriented environments.
+VII-C Expanding Sustainability Evaluation Across Programming Languages
 Another important direction for future research is the exploration of sustainability-aware code generation across a broader range of programming languages.
@@
 Although these languages are widely used in both research and industry, they do not fully capture the diversity of real-world software systems and deployment environments.
-This direction is particularly important from a sustainability perspective, as energy efficiency can vary significantly across programming languages and execution environments.
+This direction is particularly important from sustainability perspective, as energy efficiency can vary significantly across programming languages and execution environments.
 Different languages introduce different abstractions, memory management mechanisms, and runtime behaviors, all of which influence energy consumption.
@@
 Expanding sustainability evaluation across diverse programming ecosystems would provide a more comprehensive and realistic understanding of energy-aware code generation in real-world deployment scenarios.
-V-D Toward Sustainability-Aware LLM Optimization
+VII-D Toward Sustainability-Aware LLM Optimization
 Another important direction for future research is the development of sustainability-aware fine-tuning and model adaptation techniques.
-Our review shows that only one study [60] applies fine-tuning or reinforcement learning–based adaptation with an explicit focus on improving energy efficiency.
+Our review shows that no study used model adaptation techniques, e.g., fine-tuning with an explicit focus on improving energy efficiency.
 This suggests that most existing work measures sustainability at the output level without modifying the underlying model to generate more energy-efficient code.
@@
 Moreover, parameter-efficient adaptation methods can further reduce training and deployment cost, making sustainability-aware code generation more practical.
-By embedding sustainability objectives directly into model training and inference processes, future research can enable LLMs to proactively generate optimized, energy-efficient code rather than relying solely on post-hoc energy evaluation.
-V-E Limited Focus on Small Models
+By embedding sustainability objectives directly into model training and inference processes, future research can enable LLMs to proactively generate energy-efficient code rather than relying solely on post-hoc energy evaluation.
+VII-E Limited Focus on Small Models
 Another important observation from our review is the limited focus on small language models in sustainability-aware code generation research.
@@
 A deeper understanding of the accuracy–energy trade-off is essential for designing sustainable AI systems that achieve an effective balance between performance, resource utilization, and environmental impact.
-V-F Need for Sustainability-Oriented Benchmarks
+VII-F Need for Sustainability-Oriented Benchmarks
 Another important observation from this study is that none of the selected works introduce a benchmark specifically designed to evaluate sustainability.
@@
 These benchmarks should include tasks and evaluation methods that measure energy efficiency, carbon impact, resource usage, and the ability of models to produce energy-efficient solutions.
-Creating such benchmarks would improve comparability across studies and support more consistent and rigorous evaluation of sustainable code generation approaches.
-V-G Energy Measurement and Hardware Evaluation Gaps
+Creating such benchmarks would improve comparability across studies and support more rigorous evaluation of sustainable code generation approaches.
+VII-G Energy Measurement and Hardware Evaluation Gaps
 An important research gap identified in this review relates to how energy consumption is measured and evaluated across hardware platforms.
@@
 Such efforts are essential to ensure that conclusions about energy-efficient code generation are reliable, generalizable, and applicable to real-world deployment scenarios.
-VI Limitations and Threats to Validity
+VIII Limitations and Threats to Validity
 We finally offer a reflection on the validity threats that are relevant to our study and describe steps we have taken to alleviate those threats.
 We specifically focus on construct validity, internal validity, and external validity.
-VI-A Construct Validity
+VIII-A Construct Validity
 Construct validity concerns whether the key concepts of this review, sustainable code generation and LLM-based code generation, have been correctly defined and applied during the study selection and analysis process.
@@
 Although these steps do not eliminate the risk of missing or misinterpreting relevant studies, they help ensure a systematic and consistent identification of the literature included in this review.
-VI-B Internal Validity
+VIII-B Internal Validity
 Internal validity concerns whether the conclusions of this review are properly supported by the evidence presented in the selected studies.
@@
 If details were not clearly reported, we did not make assumptions.
-For instance, when a study used GitHub Copilot without specifying the underlying LLM, it was categorized as “not specified.” Although these steps cannot completely remove all bias, they help maintain a systematic, transparent, and consistent review process.
-VI-C External Validity
+For instance, when a study used GitHub Copilot without specifying the underlying LLM, it was categorized as “not specified.” Although these steps cannot completely remove all bias, they help maintain a transparent, and consistent review process.
+VIII-C External Validity
 External validity concerns whether the selected studies represent the existing body of knowledge in sustainable code generation using LLMs.
@@
 Despite these limitations, the search string, inclusion and exclusion criteria, and data extraction forms presented in this study provide a transparent framework that can be reused or extended by future researchers to replicate or expand this review.
-VII Conclusions
+IX Conclusions
 LLMs are increasingly being integrated into software engineering workflows to assist developers with tasks such as code generation, completion, translation, and bug fixing.
@@
 We conduct a systematic literature review to examine whether LLMs can generate sustainable code.
-We follow a structured process that includes comprehensive database searches and snowballing method, and applying clear inclusion and exclusion criteria in multiple screening stages.
+We follow a structured process that includes comprehensive database searches and a snowballing method, and applying clear inclusion and exclusion criteria in multiple screening stages.
 For the selected studies, we extract data using predefined forms to ensure consistency and accuracy.